### v3

# Kac-Moody Algebras on Soft Group Manifolds

Abstract
Within the so-called group geometric approach to (super)gravity and (super)string theories, any compact Lie group manifold $G_{c}$ can be smoothly deformed into a group manifold $G_{c}^{\mu}$ (locally diffeomorphic to $G_{c}$ itself), which is ‘soft’, namely, based on a non-left-invariant, intrinsic one-form Vielbein $\mu$, which violates the Maurer-Cartan equations and consequently has a non-vanishing associated curvature two-form. Within the framework based on the above deformation (‘softening’), we show how to construct an infinite-dimensional (infinite-rank), generalized Kac-Moody (KM) algebra associated to $G_{c}^{\mu}$, starting from the generalized KM algebras associated to $G_{c}$. As an application, we consider KM algebras associated to deformed manifolds such as the ‘soft’ circle, the ‘soft’ two-sphere and the ‘soft’ three-sphere. While the generalized KM algebra associated to the deformed circle is trivially isomorphic to its undeformed analogue, and hence not new, the ‘softening’ of the two- and three- sphere includes squashed manifolds (and in particular, the so-called Berger three-sphere) and yields to non-trivial results.
rutwig@ucm.es
jazzphyzz@gmail.com
Michel.Rausch@iphc.cnrs.fr
[…]
1 Instituto de Matemática Interdisciplinar and Dp.to de Geometría y Topología,
UCM, E-28040 Madrid, Spain
2 Elementar, Divisione Ricerca e Sviluppo, Galleria Enzo Tortora 21, I-10121 Torino, Italy
3 Dipartimento di Management ‘Valter Cantino’, Universitá di Torino,
Corso Unione Sovietica 218bis, I-10134 Torino, Italy
4 Department of Physics, Astronomy and Mathematics, University of Hertfordshire,
Hatfield, Hertfordshire, AL10 9AB, UK
5 Université de Strasbourg, CNRS, IPHC UMR7178, F-67037 Strasbourg Cedex, France
1 Introduction
Kac-Moody (KM) algebras, a fascinating generalization of finite-dimensional Lie algebras, have been, and still are, a cornerstone of modern Theoretical Physics and Mathematical Physics. Since the early 80’s, their emergence in various domains, from string theory and conformal field theory (CFT) to Yang-Mills theory, integrable systems and quantum groups, underscores their fundamental role in understanding the symmetry and structure of physical systems (see for instance [1]).
The affine extension of the loop algebra of smooth maps from the unit circle $\mathbb{S}^{1}$ into a simple Lie group allows to construct KM (or, better, affine) Lie algebras [2]-[5], as an alternative to their axiomatic construction (which is instead essentially based on the relaxation of the property of positive definiteness of the corresponding Cartan matrix). Generally, KM algebras are classified into three types: I) finite type, corresponding to finite-dimensional semisimple Lie algebras; II) affine type, related to loop algebras and key in two-dimensional CFT; III) indefinite type, the least understood class, potentially linked to hyperbolic symmetries. This latter type plays an important role in cosmology, as for instance in the so-called Belinskii-Khalatnikov-Lifshitz (BKL) scenario, in which the chaotic oscillations of the metric of space-time near a singularity resemble a billiard motion governed by hyperbolic KM algebras ([6]; for applications in supergravity, see e.g. [7] and Refs. therein).
In two-dimensional CFT’s, affine KM algebras describe the local symmetry. These algebras provide a rich framework for studying primary fields and their correlation functions, and also discriminate among the various universality classes in the statistical physics of critical phenomena. On the other hand, in string theory the constraints of the string’s world-sheet are governed by the Virasoro algebra, which is a central extension of the Witt algebra (i.e., of the algebra of polynomial vector fields on $\mathbb{S}^{1}$), and is intimately connected to affine Kac-Moody algebras via the Sugawara construction (see e.g. [8, 9]).
Along the years, a number of extensions and generalizations of KM algebras has been introduced and applied in a variety of contexts. Just to name a few, it is here worth mentioning quasi-simple Lie algebras [10], the generalization of KM algebras introduced in [11], and the extended Borcherds KM algebras [12]: these latter allow for imaginary simple roots, and find applications in supergravity (see e.g. [13]) as well as within the Monstrous Moonshine [14]-[16].
The Witt and the Virasoro algebras, and more generally the KM algebras, are intimately related to the compact one-dimensional manifold $\mathbb{S}^{1}$. This fact hints for the following tantalizing question: do other more general, infinite-dimensional KM algebras, related to higher-dimensional compact manifolds, exist? This is ultimately motivated also by higher-dimensional physical theories, in which harmonic expansions à la Kaluza-Klein play a crucial role (see e.g. [17]-[19], with the latter reference being motivated in the supergravity context).
The answer to the above question is positive, and was formulated in [20] (see also [21]-[24]), in which a broad class of ‘generalized KM algebras’, based on spaces of differentiable maps from compact manifolds $\mathcal{M}$ to compact Lie groups $G$, was introduced and investigated, then restricting $\mathcal{M}$ to be a compact Lie group manifold itself ($\mathcal{M}=G_{c}$) or a coset thereof ($\mathcal{M}=G_{c}/H$, where $H$ is a closed subgroup of $G_{c}$); such a restriction implied major simplifications in the treatment, because the harmonic functions on $\mathcal{M}$ could then be classified in terms of the representation theory of $G_{c}$ itself, and subsequently the Peter-Weyl theorem could be applied. In the past, such generalized KM algebras have been considered by various authors for specific manifolds, such as the two-sphere [25] or the $n$-tori [10, 25, 26, 27]; it is also here worth mentioning that, within the formulation of an extension of general relativity to closed string field theory, in [28] the possibility of generalizing KM algebras by replacing $\mathbb{S}^{1}$ with a compact coset $G_{c}/H$ was put forward.
All the aforementioned generalization of KM algebras lie outside Kac’s classification of KM algebras [3]; nevertheless, they can ultimately be regarded as generalizations of affine Lie algebras, admitting roots but not simple roots (and thus, in many cases, they do not admit a Cartan matrix at all). Moreover, unlike the usual KM algebras (whose generators are only iteratively known, level by level, by means of the Chevalley-Serre relations), the generators of such generalized KM algebras can all be constructed explicitly.
The further, somewhat natural, generalization of such algebras consisting in the relaxation of the condition of compactness of $\mathcal{M}$, has been considered in [29] (see also [30]), by focusing on the toy model provided by $\mathcal{M}=SL(2,\mathbb{R})$ and on the related symmetric coset $SL(2,\mathbb{R})/U(1)$. The lack of compactness makes the harmonic analysis on such manifolds highly non-trivial, and one must resort to different methods (with respect to the ones exploited in [20]) in order to extract suitable bases of the corresponding Hilbert spaces; in particular, the treatment given in [29] has a twofold nature: on the one hand, the Peter-Weyl theorem had to be superseded by the Plancherel theorem (generally displaying discrete and continuous series of representations of $SL(2,\mathbb{R})$ itself), while on the other hand a Hilbert basis on the space of square-integrable functions $L^{2}(SL(2,\mathbb{R}))$ was identified. The appearance of $SL(2,\mathbb{R})/U(1)$ as the target space of one complex scalar field in the bosonic sector of some Maxwell-Einstein supergravity theories in $D=3+1$ space-time dimensions would hint at an application of the resulting generalized KM algebra $\widehat{\mathfrak{g}}(SL(2,\mathbb{R})/U(1))$ in the context of supergravity (as preliminarily discussed in [29]).
In this paper, we present a further generalization of the above class of infinite-dimensional (infinite-rank) generalized KM algebras, considering the compact group manifold $G_{c}$ to be ‘deformed’ into a so-called ‘soft’ group manifold $G_{c}^{\mu}$, locally diffeomorphic to $G_{c}$ itself (the same can be done for cosets of $G_{c}$, as well). Usually, within the so-called group-geometric approach to (super)gravity and superstring theories, group manifolds are ‘softened’ in order to become domains of definition of gravitational dynamical fields, in such a way to regard $G_{c}^{\mu}$ as a vacuum configuration of a gravitational theory. Group geometry provides a natural and unified formulation of gravity and gauge theories, such that the invariances of both are interpreted as diffeomorphisms on a suitable group manifold. This geometrical framework provides a systematic algorithm for the gauging of Lie algebras and the construction of (super)gravity and (super)strings Lagrangians, and was extensively developed by the research group led by Tullio Regge, and later by his disciples Riccardo D’Auria and Pietro Fré, in Turin, starting more than fifty years ago; some sketchy presentation of the main facts will be given in Sec. 2.2, but for more details we address the reader to the lectures [31], as well as to the the first of the three books on supergravity and string theories written by Castellani, D’Auria and Fré [32] (see in particular Sec. I.3 therein). The ‘softening’ essentially amounts to deforming the original, rigid structure of the group manifold $G_{c}$, whose left- or right- invariant vector fields and one-forms have (in a given chart) a fixed coordinate dependence, and whose Riemannian geometry is (locally) fixed in terms of the structure constants of (the Lie algebra $\mathfrak{g}_{c}$ of) $G_{c}$ itself.
In other words, in this paper we consider the compact group manifold $G_{c}$ be deformed (‘softened’) into the ‘soft’ compact group manifold $G_{c}^{\mu}$, which is then potentially able to describe non-trivial physical configurations. This should be regarded as a crucial step towards the application of the generalized KM algebras under consideration in (super)gravity and (super)string theories. As explicit examples, we will consider the deformations (‘softenings’) of the circle (one-sphere) $\mathbb{S}^{1}=SO(2)\simeq U(1)$, of the two-sphere $\mathbb{S}^{2}=SO(3)/SO(2)\simeq SU(2)/U(1)=\mathbb{C}P^{1}$, and of the three-sphere $\mathbb{S}^{3}=SO(4)/SO(3)\simeq\left(SU(2)\times SU(2)\right)/SU(2)\simeq SU(2)$, respectively into the deformed circle $\mathbb{S}_{F}^{1}$, deformed two-sphere $\mathbb{S}_{F}^{2}$ and deformed three-sphere $\mathbb{S}_{F}^{3}$, all enjoying remarkable physical applications. As we will see, while the deformation of $\mathbb{S}^{1}$ is actually immaterial, the ‘softened’ manifolds $\mathbb{S}_{F}^{2}$ and $\mathbb{S}_{F}^{3}$ include, as special cases, the so-called squashed two-sphere $\widetilde{\mathbb{S}}^{2}$ and squashed three-sphere (also named Berger three-sphere) $\widetilde{\mathbb{S}}^{3}$.
The plan of the paper is as follows: Sec. 2 contains the general results described above, and it is split into three subsections. In Sec. 2.1, we recall the construction of a generalized KM algebra associated to a compact group manifold $G_{c}$ (or to a coset thereof). Then, in Sec. 2.2, we recall the basic facts of the deformation (‘softening’) procedure yielding from $G_{c}$ to the ‘soft’ group manifold $G_{c}^{\mu}$. Subsequently, in Sec. 2.3 we associate a generalized KM algebra to $G_{c}^{\mu}$, starting from the generalized KM algebras associated to $G_{c}$. Sec. 3 presents the most elementary example of such a procedure, pertaining to the circle $\mathbb{S}^{1}$, in which case the generalized KM algebra associated to the deformed circle $\mathbb{S}_{F}^{1}$ is trivially isomorphic to its undeformed analogue; this is not surprising, since the topological classification of one-dimensional closed curves shows that all such curves are topologically equivalent to $\mathbb{S}^{1}$. Then, Sec. 4 presents a detailed treatment of the three-sphere $\mathbb{S}^{3}$, whose ‘softened’ version includes the so-called squashed three-sphere $\widetilde{\mathbb{S}}^{3}$. Finally, by removing the dependence on one angle, Sec. 5 deals with the two-sphere $\mathbb{S}^{2}$ and its ‘softening’ $\mathbb{S}_{F}^{2}$, which includes the so-called squashed two-sphere $\widetilde{\mathbb{S}}^{2}$. Some conclusive remarks are made in Sec. 6, and the paper is concluded by an Appendix, which recalls some basic facts on the Maurer-Cartan one-forms of $SU(2)$.
2 Kac-Moody (KM) algebras on ‘soft’ Lie group manifolds
Let $G_{c}$ be a compact Lie group. In this section we recall the salient steps for the construction of a Kac-Moody (KM) algebra associated to $G_{c}$ (for more details see [20]). In a second part, we turn to the construction of a KM algebra on a deformation (more specifically, a ‘softening’, see below) of $G_{c}$.
2.1 A KM algebra on $G_{c}$
Let $\mathfrak{g}=\{T_{1},\cdots,T_{d}\}$ be a $d-$dimensional simple Lie algebra (complex or real) with Lie brackets
$\big{[}T_{a},T_{b}\big{]}=if_{ab}{}^{c}T_{c},$ (2.1)
and Killing form
$\big{<}T_{a},T_{b}\big{>}_{0}=\eta_{ab}=\text{tr}(\text{ad}(T_{a})\text{ad}(T_{b}))\ .$
Further, let $G_{c}$ be a compact Lie group. Hence, $G_{c}$ is a compact group manifold, that we assume to be of dimension $n$. Let $m^{A}=(\varphi^{i},\theta^{r})$ with $i=1,\cdots,p;r=1,\cdots,q$ and $n=p+q$ be a parameterisation of $G_{c}$; this split of $m^{A}$ is such that the matrix elements are periodic in the $\varphi^{i}$’s and not periodic in the $\theta^{r}$’s [^1] For instance, the spherical coordinates’ parametrization of the (unit) $\mathbb{S}^{2}$ has $\varphi\in[0,2\pi)$ and $\cos\theta\in[-1,1].$. Within this parameterisation, a generic element (connected to the identity) of $G_{c}$ is given by
$g(m)=e^{im^{A}J_{A}}\ ,$
where $J_{A},A=1,\cdots,n$ are the generators of $\mathfrak{g}_{c}$, the Lie algebra of $G_{c}$, not to be confused with the generators of $\mathfrak{g}$. Then, denote the coordinates of a group element (in a local coordinate chart) by
$g(m)^{M}\equiv m^{M}\ ,\ \ M=1,\cdots,n.$
The indices $A,B,\cdots$ are tangent space indices, i.e., flat indices, whilst the indices $M,N,\cdots$ are world indices, i.e., curved indices. The Vierbein one-form associated to the aforementioned parameterisation reduces to
$e(m)=g(m)^{-1}\;\mathrm{d}g(m)\ ,$
and satisfies the Maurer-Cartan equation
$\text{d}e+e\wedge e=0\ .$ (2.2)
The components of Vielbein are
$e^{A}(m)=e_{M}{}^{A}(m)\;\text{d}m^{M}\ .$
Therefore the metric tensor on $G_{c}$ is
$g_{MN}=e_{M}{}^{A}(m)e_{N}{}^{B}(m)\delta_{AB}\ .$ (2.3)
We consider now the Hilbert space $L^{2}(G_{c})$, i.e., the set of square integrable functions on $G_{c}$, endowed with the scalar product:
$(f,g)=\frac{1}{V}\int_{G_{c}}\sqrt{g}\;\text{d}\varphi^{p}\text{d}\theta^{q}\overline{f(\varphi,\theta)}g(\varphi,\theta),$ (2.4)
where $g=\det(g_{MN})$ and $V$ is the volume of $G_{c}$. Since $G_{c}$ is compact, its unitary representations are finite dimensional and it turns out that, once correctly normalized, the set of all matrix elements constitutes an orthonormal Hilbert basis of $L^{2}(G_{c})$. This is the Peter–Weyl Theorem [33]. Introduce $\mathcal{I}$ the minimal countable set of labels required to identify the states unambiguously. With this notation, one can write the matrix elements as $\rho_{I}(\varphi,\theta)$ ([20], see also below) and the Hilbert basis of $L^{2}(G_{c})$ is given by
=$\mathcal{B}=\Big{\{}\rho_{I}(\varphi,\theta)\ ,\ \ I\in\mathcal{I}\Big{\}}\ ,$ (2.5)
and the matrix elements are orthonormal with respect to the scalar product (2.4)
$(\rho_{I},\rho_{J})=\delta_{IJ}\ .$
Since the product of two different elements of the basis $\mathcal{B}$ is square integrable, we have the following decomposition
$\rho_{I}(\varphi,\theta)\rho_{J}(\varphi,\theta)=c_{IJ}{}^{K}\rho_{K}(\varphi,\theta)$ (2.6)
where the coefficients $c_{IJ}{}^{K}=c_{JI}{}^{K}$ are the Clebsch-Gordan coefficients of $G_{c}$ [20].
The parameterisation $m^{A}=(\varphi^{i},\theta^{r})$ leads naturally to a differential realization of the Lie algebra $\mathfrak{g}_{c}$ for the generators of the left and right action. Let $L_{A}$ (resp. $R_{A}$) $A=1,\cdots n$, be the generators of the left (resp. right) action satisfying
$\big{[}L_{A},L_{B}\big{]}=iC_{AB}{}^{C}L_{C}\ ,\ \ \big{[}R_{A},R_{B}\big{]}=iC_{AB}{}^{C}R_{C}\ ,\ \ \big{[}L_{A},R_{B}\big{]}=0\ ,$
where $C_{AB}{}^{C}$ are the structure constants of $\mathfrak{g}_{c}$. Recall how one can construct in an explicit way the generators $L_{A}$ and $R_{A}$. Let $\mathcal{M}$ be a Riemannian manifold. Let $m^{M},M=1,\cdots,\dim\mathcal{M}$ be a parameterisation of $\mathcal{M}$. The manifold $\mathcal{M}$ is endowed with a metric $g$ and let $\nabla_{M}$ be the corresponding covariant derivative [34]. Consider a coordinate transformation $m^{M}\to m^{M}+\epsilon\xi^{M}$ with $\epsilon\sim 0$. This transformation is an isometry, i.e., it preserves the metric if we have the Killing equation:
$\nabla_{M}\xi_{N}+\nabla_{N}\xi_{M}=0\ ,$ (2.7)
where $\xi_{M}=g_{MN}\xi^{N}$. For any solution of the Killing equation $\xi^{M}_{A}(m)$, the differential operator
$J_{A}=-i\xi_{A}^{M}(m)\;\frac{\partial}{\partial m^{M}}$
leaves the metric invariant. When the manifold corresponds to a group manifold, the generators $L_{A},R_{A}$ (corresponding of the left and right action of $G_{c}$ onto itself) are solutions of the Killing equation (2.7).
Furthermore, recall that if $\omega$ is a one-form under the isometry $J_{A}$, we have
$\delta_{A}\omega_{M}=-\Big{(}\xi_{A}^{N}(m)\frac{\partial\omega_{M}}{\partial m^{N}}+\omega_{N}\frac{\partial\xi_{A}^{N}(m)}{\partial m^{M}}\Big{)}\ .$
Thus $\omega$ is invariant under the action of $J_{A}$ if we have
$\xi_{A}^{N}(m)\frac{\partial\omega_{M}}{\partial m^{N}}+\omega_{N}\frac{\partial\xi_{A}^{N}(m)}{\partial m^{M}}=0\ .$ (2.8)
Let $I=(L,Q,R)$ with $L=(\ell_{1},\cdots,\ell_{\frac{1}{2}(n-r_{c})}),R=(r_{1},\cdots,r_{\frac{1}{2}(n-r_{c})})$ and $Q=(q_{q},\cdots,q_{r_{c}})$, where $r_{c}$ is the rank of $\mathfrak{g}_{c}$. Indeed, a unitary representation is specified by the $r_{c}$ eigenvalues of the primitive Casimir operators, and any vector of a given representation is unambiguously specified by the eigenvalues of $\frac{1}{2}(n-r_{c})$ commuting Hermitean operators [20]. Thus $L,Q,R$ constitute the minimal set of indices to specify any vector in the basis $\mathcal{B}$ (2.5) (see [35, 36, 37] for details). The action of $L_{A},R_{A}$ reads
$L_{A}\rho_{LQR}(\varphi,\theta)$ $=$ $C^{Q}_{AL}{}^{L^{\prime}}\rho_{L^{\prime}QR}(\varphi,\theta)$
$R_{A}\rho_{LQR}(\varphi,\theta)$ $=$ $C^{Q}_{AR}{}^{R^{\prime}}\rho_{LQR^{\prime}}(\varphi,\theta)$
where $C^{Q}_{AL}{}^{L^{\prime}}$ (resp. $C^{Q}_{AR}{}^{R^{\prime}}$) are the matrix elements of the left (right) action for the representation specified by $Q$.
We next define a space of smooth mappings from $G_{c}$ into $\mathfrak{g}$ as
$\mathfrak{g}(G_{c})=\Big{\{}T_{aI}=T_{a}\rho_{I}(\varphi,\theta)\ ,\ \ a=1,\cdots,d\ ,\ \ I\in\mathcal{I}\Big{\}}\ ,$ (2.9)
which inherits the structure of a Lie algebra
$\big{[}T_{aI},T_{bJ}\big{]}=if_{ab}{}^{c}c_{IJ}{}^{K}T_{cK}.$ (2.10)
This, in particular means that the Lie brackets of $\mathfrak{g}(G_{c})$ are obtained in terms of the structure constants $f_{ab}{}^{c}$ of $\mathfrak{g}$ and the Clebsch-Gordan coefficients $c_{IJ}{}^{K}$ of $G_{c}$.
The last step in the construction of the KM algebra associated to $G_{c}$ is to introduce Hermitean operators and central charges in duality. Of course, as the variables $\varphi^{i}$ are periodic, the operators $-i\partial_{\varphi^{i}}$ are Hermitean. However, additional Hermitian operators can be considered. For instance, the $2r_{c}$ set of commuting operators obtained from $L_{A},R_{B}$ and corresponding to the Cartan subalgebra of $\mathfrak{g}_{c}$ constitutes a set of $2n_{c}$ Hermitean commuting operators. In some case, however, this number can be larger. The existence of these additional commuting operators has been analysed in [20]. Denote $D_{i},i=2r_{c}\leq 1,\cdots,r\leq n$ the set of commuting Hermitean operators ($r=n$ only in the case of the $n-$tori $\mathbb{T}_{n}=U(1)^{n}$). The elements $T_{aI}$ are eigenfunctions of $D_{i}$:
$\big{[}D_{i},T_{aI}\big{]}$ $=$ $I(i)T_{aI}\ .$ (2.11)
There exists in principle an infinite number of central charges, hence we limit ourselves to $r$ central charges $k_{1},\cdots,k_{r}$ in duality with the Hermitean operators, and given by the two-cocycle (see also [38, 39]):
$\omega_{i}(T_{aI},T_{bJ})$ $=$ $\frac{k_{i}}{V}\int_{G_{c}}\sqrt{g}\text{d}^{p}\varphi\text{d}^{q}\theta\Big{<}T_{aI},D_{i}T_{bJ}\Big{>}_{0}$ (2.12)
$=$ $k_{i}J(i)\eta_{ab}\eta_{IJ}$
where $\eta_{IJ}=\pm\delta_{IJ}$ (see [20] for more details).
The KM algebra associated to $G_{c}$, denoted $\widetilde{\mathfrak{g}}(G_{c})$, is a central extension of the algebra ${\mathfrak{g}}(G_{c})$. We denote $\mathcal{T}_{aI}$ the generators in $\widetilde{\mathfrak{g}}(G_{c})$ corresponding to the generators $T_{aI}=T_{a}\rho_{I}(\varphi,\theta)$ in ${\mathfrak{g}}(G_{c})$. Thus $\widetilde{\mathfrak{g}}(G_{c})=\big{\{}\mathcal{T}_{aI},a=1\cdots,\dim\mathfrak{g},I\in\mathcal{I},D_{i},k_{i},i=1,\cdots r\}$. From (2.10), (2.11) and (2.12) the Lie brackets are
$\big{[}\mathcal{T}_{aI},\mathcal{T}_{bJ}\big{]}$ $=$ $if_{ab}{}^{c}c_{IJ}{}^{K}\mathcal{T}_{cK}+\eta_{ab}\eta_{IJ}\sum\limits_{i=1}^{r}k_{i}J(i)$
$\big{[}D_{i},\mathcal{T}_{aI}\big{]}$ $=$ $I(i)\mathcal{T}_{aI}\ .$ (2.13)
All the results of this section extend naturally to the coset space $G/H$, where $H$ is a subgroup of $G$ [20].
2.2 ‘Softening’ of (compact) Lie group manifolds
We consider a smooth, ‘softening’ deformation $G_{c}^{\mu}$ of the Lie group $G_{c}$, locally diffeomorphic to $G_{c}$ itself [^2] Here, we are not going to deal with th general theory of the ‘softening’ of compact cosets $G_{c}/H$ (involving the so-called ‘horizontality condition’ for the curvatures), addressing the reader to Sec. I.3.7 of [32] for a treatment and a list of Refs.. (see e.g. [31] for a review, and a list of Refs.). We assume that the Vielbein $\mu$ is an intrinsic one-form (valued in the Lie algebra $\mathfrak{g}_{c}$ of $G_{c}$)
$\mu^{A}(m)=\mu_{A}{}^{M}(m)\;\text{d}m^{M}$
i.e., it is not a Maurer-Cartan one-form (namely, it does not satisfy (2.2)). In other words, $\mu$ is not left-invariant (i.e., it is a ‘soft’, intrinsic one-form), and it does not satisfy the Maurer-Cartan equation, but rather it holds that
$\text{d}\mu+\mu\wedge\mu=R,$ (2.14)
where $R$ is the curvature two-form of $\mu$. It is in this sense that we consider that the ‘soft’ group manifold $G_{c}^{\mu}$ is a deformation [^3] In the case $ISO(1,3)/SO(1,3)$, namely the coset of the Poincaré group by the Lorentz group, the analogue of (2.3) in this case leads to the flat Minkowski space-time, whilst the intrinsic definition of the Vielbein above leads to a Riemannian space with curvature (see e.g. Sec. I.3 of [32]). of $G_{c}$.
Then, $\mu$ span a basis of the cotangent plane of $G_{c}^{\mu}$, and one can define the metric tensor
$g_{MN}^{\mu}(m)=\mu_{M}{}^{A}(m)\mu_{N}{}^{B}(m)\delta_{AB}.$ (2.15)
Taking the exterior derivative of both sides of Eq. (2.14), one obtains the Bianchi identity for the curvature of $\mu$,
$dR+2R\wedge\mu=0\Leftrightarrow\nabla R=0,$
where the covariant derivative operator $\nabla$ on $G_{c}^{\mu}$ has been introduced.
We will further assume that the manifold $G_{c}^{\mu}$ has the same parameterisation $m^{M}=(\varphi^{i},\theta^{r})$, the only difference between $G_{c}$ and $G_{c}^{\mu}$ being at the level of the metric tensor ((2.3) or (2.15) respectively). The corresponding scalar product on $G_{c}^{\mu}$ is then given by
$(f,g)_{\mu}=\frac{1}{V}\int_{G_{c}}\sqrt{g^{\mu}}\;\text{d}\varphi^{p}\text{d}\theta^{q}\overline{f(\varphi,\theta)}g(\varphi,\theta),$ (2.16)
where $g^{\mu}=\det(g_{MN}^{\mu})$. Notice that, since the parameterisation of $G_{c}^{\mu}$ and $G_{c}$ is the same, the limits of integration are again $G_{c}$ in this case.
Two brief remarks are in order (for further elucidation, see e.g. [31]).
  1. 1.
    The definition of ‘soft’ one-forms $\mu$ and of the associated curvature two-form $R$ is the same as in Yang-Mills theory, with the crucial difference that in the present case, the Vielbein one-form $\mu$ is defined on $G_{c}^{\mu}$, which does not have an apriori fiber bundle structure.
  2. 2.
    We have introduced the ‘soft’ forms starting from the dual covariant formulation of the Lie algebra $\mathfrak{g}_{c}$ of $G_{c}$, namely from the Maurer-Cartan equation (2.2). Of course, the same can be done in the contravariant language of vector fields.
2.3 A KM algebra on $G_{c}^{\mu}$
Let $L^{2}(G_{c}^{\mu})$ be the set of square integrable functions on $G^{\mu}_{c}$ endowed with the scalar product (2.16). Following Mackey [40], we can easily construct a complete set of orthonormal functions on $L^{2}(G_{c}^{\mu})$ (see also [29]). Recall the main points which enable to associate a Hilbert basis on a manifold $\mathcal{M}$ with integration measure $\text{d}\beta$ starting from a Hilbert basis on the same manifold $\mathcal{M}$ with integration measure $\text{d}\alpha$ ([40], p. 100). These two integration measures endow $\mathcal{M}$ with two different scalar products
$\begin{array}[]{ll}(\mathcal{M},\text{d}\alpha):&(f,g)_{\alpha}=\int_{\mathcal{M}}\text{d}\alpha\;\overline{f(m)}g(m)\\[5.69046pt] (\mathcal{M},\text{d}\beta):&(f,g)_{\beta}=\int_{\mathcal{M}}\text{d}\beta\;\overline{f(m)}g(m)\end{array}$ (2.19)
with $m\in\mathcal{M}$. We assume further that there exists a mapping $T_{\beta\alpha}$:
$T_{\beta\alpha}:L^{2}(\mathcal{M},\text{d}\beta)\to L^{2}(\mathcal{M},\text{d}\alpha)\ ,$
such that
$\int_{\mathcal{M}}\text{d}\alpha=\int_{\mathcal{M}}\text{d}\beta T_{\beta\alpha}.$
For instance, for an $n$-dimensional Riemannian manifold $\mathcal{M}$ parameterized by $m_{1},\cdots,m_{n}$ with metric $g_{\alpha}$ (resp. $g_{\beta}$) we have $\text{d}\alpha=\sqrt{\big{|}\det g_{\alpha}\big{|}}\text{d}^{n}m$ (resp. $\text{d}\beta=\sqrt{\big{|}\det g_{\beta}\big{|}}\text{d}^{n}m$) and thus $T_{\beta\alpha}=\sqrt{\big{|}\det g_{\alpha}\big{|}/\big{|}\det g_{\beta}\big{|}}$. This means that if $\{f^{\beta}_{i},i\in\mathbb{N}\}$ is a Hilbert basis of $L^{2}(\mathcal{M},\text{d}\beta)$, then $\{f^{\alpha}_{i}=\frac{f^{\beta}_{i}}{\sqrt{T_{\beta\alpha}}},i\in\mathbb{N}\}$ is a Hilbert basis for $L^{2}(\mathcal{M},\text{d}\alpha)$, and we obviously have
$(f^{\beta}_{i},f^{\beta}_{j})_{\beta}=\delta_{ij}\ \ \Longleftrightarrow\ \ (f^{\alpha}_{i},f^{\alpha}_{j})_{\alpha}=\delta_{ij}$
and the map $T_{\beta\alpha}$ is unitary.
Applied to our case, the transition functions read $T^{\mu}(m)=\sqrt{\frac{g}{g^{\mu}}}$ (see (2.16)) and the orthonormal Hilbert basis of $G_{c}^{\mu}$ is
$\mathcal{B}_{\mu}=\Big{\{}\rho_{I}^{\mu}(\varphi,\theta)=\sqrt{T^{\mu}}\rho_{I}(\varphi,\theta)\ ,\ \ I\in\mathcal{I}\Big{\}},$ (2.20)
trivially satisfying the relation
$(\rho_{I}^{\mu},\rho^{\mu}_{J})_{\mu}=\delta_{IJ}\ .$
If we now introduce
$L_{A}^{\mu}=\sqrt{T^{\mu}}L_{A}\frac{1}{\sqrt{T^{\mu}}}\ ,\ \ R_{A}^{\mu}=\sqrt{T^{\mu}}R_{A}\frac{1}{\sqrt{T^{\mu}}}\ ,$
it follows at once that
$\big{[}L^{\mu}_{A},L^{\mu}_{B}\big{]}=iC_{AB}{}^{C}L^{\mu}_{C}\ ,\ \ \big{[}R^{\mu}_{A},R^{\mu}_{B}\big{]}=iC_{AB}{}^{C}R^{\mu}_{C}\ ,\ \ \big{[}L^{\mu}_{A},R^{\mu}_{B}\big{]}=0\ .$
and
$L^{\mu}_{A}\rho^{\mu}_{LQR}(\varphi,\theta)$ $=$ $C^{Q}_{AL}{}^{L^{\prime}}\rho^{\mu}_{L^{\prime}QR}(\varphi,\theta)$
$R^{\mu}_{A}\rho_{LQR}(\varphi,\theta)$ $=$ $C^{Q}_{AR}{}^{R^{\prime}}\rho^{\mu}_{LQR^{\prime}}(\varphi,\theta)$ (2.21)
Thus $\{L_{A}^{\mu},A=1,\cdots,n\}$ and $\{R_{A}^{\mu},A=1,\cdots,n\}$ generate the Lie algebra $\mathfrak{g}_{c}$ and $\rho^{\mu}_{LQR}$ are the corresponding matrix elements of $G_{c}$ (but not of $G_{c}^{\mu}$, which is not a group).
To define the analogue of (2.9) for the deformed, ‘soft’ group manifold $G_{c}^{\mu}$, we need more. Indeed, the products $\rho_{I}^{\mu}\rho_{J}^{\mu}$ must be square integrable for any $I,J.$ We now show that if for any $I,J$ the function $\rho_{I}^{\mu}\rho_{I}^{\mu}\in L^{2}(G_{c}^{\mu})$, then $\sqrt{T^{\mu}}\in L^{2}(G_{c})$. Indeed, for the trivial representation of $G_{c}$ it holds that $\rho_{0}(\varphi,\theta)=1$, thus by hypothesis $(\rho_{0}^{I})^{2}=T^{\mu}\in L^{2}(G_{c}^{\mu})$. Now $(\rho_{0}^{I})^{2}\in L^{2}(G_{c}^{\mu})$ leads to $\sqrt{T^{\mu}}\in L^{2}(G_{c})$ because
$\frac{1}{V}\int_{G_{c}}\sqrt{g^{\mu}}\;\text{d}\varphi^{p}\text{d}\theta^{q}(T^{\mu}(\varphi,\theta))^{2}=\frac{1}{V}\int_{G_{c}}\sqrt{g}\;\text{d}\varphi^{p}\text{d}\theta^{q}{T^{\mu}(\varphi,\theta)}$ (2.22)
Consequently, $\sqrt{T^{\mu}}$ is a square integrable function of $L^{2}(G_{c})$ (we already know that indeed $\sqrt{T^{\mu}}$ is a square integrable function of $L^{2}(G_{c}^{\mu})$) and we have
$\sqrt{T^{\mu}}=C^{I}\rho_{I}(\varphi,\theta)$ (2.23)
Since now by hypothesis the product $\rho^{\mu}_{I}\rho^{\mu}_{J}$ belongs to $L^{2}(G_{c}^{\mu})$:
$\rho^{\mu}_{I}(\varphi,\theta)\rho^{\mu}_{J}(\varphi,\theta)=c^{\mu}_{IJ}{}^{K}\rho^{\mu}_{K}(\varphi,\theta)$ (2.24)
where, using (2.23),
$c^{\mu}_{IJ}{}^{K}\equiv C^{L}c_{IJ}^{~{}~{}M}c_{LM}^{~{}~{}K}$ (2.25)
can be regarded as the [^4] Here and below, these quotes are used to stress the slight abuse of language, due to the fact that $G_{c}^{\mu}$ is not a group. ‘Clebsch-Gordan coefficients’ of $G_{c}^{\mu}$.
Conversely, if the function $\sqrt{T^{\mu}}\in L^{2}(G_{c})$ is sufficiently smooth, then $\sqrt{T^{\mu}}\in L^{2}(G_{c})$ implies that for any $I,J$, the product $\rho^{\mu}_{I}\rho^{\mu}_{J}\in L^{2}(G_{c}^{\mu})$. Indeed, (because of (2.23) and (2.6)) we have
$\rho^{\mu}_{I}(\varphi,\theta)\rho^{\mu}_{J}(\varphi,\theta)$ $=$ $\sqrt{T^{\mu}(\varphi,\theta)}\sqrt{T^{\mu}(\varphi,\theta)}\rho_{I}(\varphi,\theta)\rho_{J}(\varphi,\theta)$
$=$ $\sqrt{T^{\mu}(\varphi,\theta)}C^{K}c_{IJ}{}^{M}\rho_{K}(\varphi,\theta)\rho_{M}(\varphi,\theta)$
$=$ $C^{K}c_{IJ}{}^{M}c_{KM}{}^{N}\rho^{\mu}_{N}(\varphi,\theta)$
$=$ $c^{\mu}_{IJ}{}^{N}\rho_{N}^{\mu}(\varphi,\theta)\ .$
As seen previously the $\rho_{I}^{\mu}$ are in the left and right representation of $G_{c}$ (see (2.3)). However, since the metric tensor is deformed by the parameter $T^{\mu}$, we have to take into account this deformation parameter when considering tensor product of representations. In particular, we define
$\rho_{I}^{\mu}(\varphi,\theta)\otimes_{\mu}\rho_{J}^{\mu}(\varphi,\theta)\equiv\frac{1}{\sqrt{T^{\mu}}}\rho_{I}^{\mu}(\varphi,\theta)\rho_{J}^{\mu}(\varphi,\theta)=\sqrt{T^{\mu}}\rho_{I}(\varphi,\theta)\rho_{J}(\varphi,\theta)=c_{IJ}{}^{K}\rho_{K}^{\mu}(\varphi,\theta)\ ,$
with $c_{IJ}{}^{K}$ the Clebsch-Gordan coefficients of $G_{c}$, and thus recovering the usual results (see (2.6)). Note that this equation is very different from (2.24). Indeed, in (2.24) we have considered the usual tensor product inherited from the group $G_{c}$, whilst in the definition above we have used the ‘deformed’ tensor product associated to $G_{c}^{\mu}$.
Thus, under these conditions,
$\mathfrak{g}(G_{c}^{\mu})=\Big{\{}T_{aI}^{\mu}=T_{a}\rho_{I}^{\mu}(\varphi,\theta)\ ,\ \ a=1,\cdots,d\ ,\ \ I\in\mathcal{I}\Big{\}}\ ,$
is a Lie algebra with Lie brackets
$\big{[}T_{aI}^{\mu},T_{bJ}^{\mu}\big{]}=if_{ab}{}^{c}c_{IJ}^{\mu}{}^{K}T_{cK}^{\mu}\ .$
The definition of Hermitian and central charges follows easily. For the Hermitean commuting operators we introduce
$D_{i}^{\mu}=\sqrt{T^{\mu}}D_{i}\frac{1}{\sqrt{T_{\mu}}}\ ,$
which obviously leads to
$\big{[}D_{i}^{\mu},T_{aI}^{\mu}\big{]}=I(i)T_{aI}^{\mu}\ .$
Similarly, since
$T_{aI}^{\mu}=T_{a}\rho_{I}^{\mu}=\sqrt{T^{\mu}}T_{a}\rho_{I}=\sqrt{T^{\mu}}T_{aI}=T_{a}C^{J}\rho_{I}\rho_{J}=T_{a}C^{J}c_{IJ}^{~{}~{}K}\rho_{K}=C^{J}c_{IJ}^{~{}~{}K}T_{aK}=P_{I}^{~{}K}T_{aK},$
where [^5] Such that the ‘Clebsch-Gordan coefficients’ of $G_{c}^{\mu}$, defined by (2.25), are given by $c^{\mu}_{IJ}{}^{K}=c_{IJ}^{~{}~{}M}P_{M}^{~{}K}$.
$P_{I}^{~{}K}\equiv C^{J}c_{IJ}^{~{}~{}K},$ (2.26)
the two-cocycles are given by
$\omega_{i}^{\mu}\left(T_{aI}^{\mu},T_{bJ}^{\mu}\right)=\omega_{i}\left(P_{I}^{~{}I^{\prime}}T_{aI^{\prime}},P_{J}^{~{}J^{\prime}}T_{bJ^{\prime}}\right)=P_{I}^{~{}I^{\prime}}P_{J}^{~{}J^{\prime}}\omega_{i}\left(T_{aI^{\prime}},T_{bJ^{\prime}}\right)=k_{i}\eta_{ab}P_{I}^{~{}I^{\prime}}P_{J}^{~{}J^{\prime}}J^{\prime}(i)\eta_{I^{\prime}J^{\prime}}.$ (2.27)
Note that $\omega_{i}^{\mu}$ is obtained by an integration over the original manifold $G_{c}$, and not over $G_{c}^{\mu}$. This in particular means that the differential operators that we have to take in duality with the two-cocycles are the original Hermitean operators $D_{i}$, and not the operators $D_{i}^{\mu}$. The fact that $\omega_{i}$ are two-cocycles ensure naturally that $\omega_{i}^{\mu}$ are also two-cocycles. Indeed, both cocycles are defined by the same integration upon $G_{c}$.
The KM algebra associated to $G^{\mu}_{c}$ is given by $\widetilde{\mathfrak{g}}(G^{\mu}_{c})=\big{\{}\mathcal{T}^{\mu}_{aI}\ ,\ \ a=1,\cdots,d\ ,I\in\mathcal{I},D_{i},k_{i},i=1,\cdots r\}$, where $\mathcal{T}_{aI}^{\mu}$ are the generators in $\widetilde{\mathfrak{g}}(G^{\mu}_{c})$ corresponding to the generators $T_{aI}^{\mu}=T_{a}\rho_{I}^{\mu}(\varphi,\theta)$ in $\mathfrak{g}(G_{c}^{\mu})$. The Lie brackets read [^6] Despite having adopted the Einstein summation convention on repeated indices, here we keep the $\sum_{i=1}^{r}$ symbol to stress the sum over all $r$ central extensions.
$\left[\mathcal{T}_{aI}^{\mu},\mathcal{T}_{bJ}^{\mu}\right]$ $=$ $if_{ab}^{~{}~{}c}c^{\mu}_{IJ}{}^{K}\mathcal{T}_{cK}^{\mu}+\sum_{i=1}^{r}\omega_{i}^{\mu}\left(\mathcal{T}_{aI}^{\mu},\mathcal{T}_{bJ}^{\mu}\right)=$
$=$ $if_{ab}^{~{}~{}c}c^{\mu}_{IJ}{}^{K}\mathcal{T}_{cK}^{\mu}+P_{I}^{~{}I^{\prime}}P_{J}^{~{}J^{\prime}}\eta_{ab}\eta_{I^{\prime}J^{\prime}}\sum_{i=1}^{r}k_{i}J^{\prime}(i);$
$\left[D_{i},\mathcal{T}_{aI}^{\mu}\right]$ $=$ $P_{I}^{~{}I^{\prime}}I^{\prime}(i)\mathcal{T}_{aI^{\prime}}^{\mu}.$ (2.28)
[…]
3 ‘Softening’ of $\mathbb{S}^{1}$ and the associated KM algebra
As a first example, let us study a KM algebra associated to the ‘softened’ deformation of the one-sphere (circle) $\mathbb{S}^{1}$. The KM algebra associated to the undeformed circle $\mathbb{S}^{1}$ is nothing but the affine Lie algebra $\widetilde{\mathfrak{g}}$.
The Hilbert space $L^{2}(\mathbb{S}^{1})$ is endowed with the natural scalar product:
$(f,g)=\frac{1}{2\pi}\int\limits_{0}^{2\pi}\text{d}\theta\;\overline{f(\theta)}g(\theta)\ ,$ (3.1)
and the Hilbert basis is
=$\mathcal{B}=\big{\{}e_{n}(\theta)=e^{in\theta}\ ,\ \ n\in\mathbb{Z}\big{\}}\ .$
Since
$e_{n}\left(\theta\right)e_{m}\left(\theta\right)=e_{n+m}\left(\theta\right),$
the Clebsch-Gordan coefficients of $\mathbb{S}^{1}$ are simply given by
$c_{mn}^{~{}~{}p}=\delta_{p,m+n}.$ (3.2)
Let $d=-i\partial_{\theta}$, and $\mathfrak{g}({\mathbb{S}^{1}})=\{T_{am}=T_{a}e_{m}(\theta)\}$ be the loop algebra associated to $\mathfrak{g}$. The central extension is associated to the two-cocycle:
$\omega(T_{am},T_{bn})=\frac{k}{2\pi}\int\limits_{0}^{2\pi}\text{d}\theta\;\Big{<}T_{am},dT_{bn}\Big{>}_{0}=kn\;\eta_{ab}\delta_{n,-m}\ .$
Thus, the affine Lie algebra (which centrally extends the loop algebra) is
$\widetilde{\mathfrak{g}}=\Big{\{}\mathcal{T}_{an},k,d,a=1,\cdots,d,m\in\mathbb{Z}\Big{\}}$
and the Lie brackets are given by
$\big{[}\mathcal{T}_{am},\mathcal{T}_{bn}\big{]}$ $=$ $if_{ab}{}^{c}\mathcal{T}_{c|m+n}+kn\eta_{ab}\delta_{m,-n}$
$\big{[}d,\mathcal{T}_{am}\big{]}$ $=$ $m\mathcal{T}_{am}\ .$ (3.3)
Let $f$ and $g$ be two functions on the circle. The algebra of vector fields (i.e., the de Witt algebra) is given by the Lie brackets
$\big{[}f(\theta)\partial_{\theta},g(\theta)\partial_{\theta}\big{]}=\Big{(}f(\theta)g^{\prime}(\theta)-g(\theta)f^{\prime}(\theta)\Big{)}\partial_{\theta}\ .$
A natural basis of the de Witt algebra is given by $\{\ell_{n},n\in\mathbb{Z}\}$ with $\ell_{n}(\theta)=ie^{in\theta}\partial_{\theta}$. Over this basis, the brackets acquire the form:
$\big{[}\ell_{m},\ell_{n}\big{]}=(m-n)\ell_{m+n}\ .$ (3.4)
The de Witt algebra admits a central charge given by the Gel’fand–Fuks cocycle [9]
$\omega(f,g)=-i\frac{c}{12}\frac{1}{2\pi}\int\limits_{0}^{2\pi}\text{d}\theta\;f(\theta)g^{\prime\prime\prime}(\theta)$ (3.5)
In particular, it follows that
$\omega(\ell_{m},\ell_{n})=\frac{c}{12}m^{3}\delta_{m+n,0}\ .$ (3.6)
Let $\mathcal{L}_{m},m\in\mathbb{Z}$ be the generators of the Virasoro algebra, i.e., the centrally extended de Witt algebra. The Lie brackets reduce to
$\big{[}\mathcal{L}_{m},\mathcal{L}_{n}\big{]}=(m-n)\mathcal{L}_{m+n}+\frac{c}{12}m^{3}\delta_{m+n,0}\ .$
A more convenient basis is given by $L_{m}=\mathcal{L}_{m}+\frac{c}{24}\delta_{n,0}$. Over this basis, we obtain the standard Lie brackets of the Virasoro algebra:
$\big{[}L_{m},L_{n}\big{]}=(m-n)L_{m+n}+\frac{c}{12}(m^{3}-m)\delta_{m+n,0}\ .$ (3.7)
The Virasoro algebra (3.7) and the affine Lie algebra (3) admit a semi-direct structure:
$\big{[}L_{m},T_{an}\big{]}=-nT_{an+m}\ ,$ (3.8)
thus the Hermitean operator $d$ of the affine Lie algebra can be identified with $-L_{0}$.
To construct a deformation of (3), following Sections 2.2 and 2.3 , we introduce the deformed scalar product on the ‘soft’ circle $\mathbb{S}_{F}^{1}$:
$(f,g)_{F}=\frac{1}{2\pi}\int\limits_{0}^{2\pi}F(\theta)\text{d}\theta\;\overline{f(\theta)}g(\theta)\ ,$ (3.9)
where $F$ is a positive function such that $1/\sqrt{F}$ belongs to $L^{2}(\mathbb{S}^{1})$. The Hilbert basis of $L^{2}(\mathbb{S}_{F}^{1})$ is
$\mathcal{B}_{F}=\Big{\{}e_{n}^{F}(\theta)=\frac{1}{\sqrt{F(\theta)}}e^{in\theta}\ ,\ \ n\in\mathbb{Z}\Big{\}}$ (3.10)
(the set of functions $\{e_{n}^{F}\}$ constitutes an orthonormal set with respect to the scalar product (3.9)). Since [^9] To be very explicit, in this Section, we do not use Einstein’s summation convention on repeated indices.
$\frac{1}{\sqrt{F\left(\theta\right)}}=\sum_{n\in\mathbb{Z}}F^{n}e_{n}\left(\theta\right)=\sum_{n\in\mathbb{Z}}F^{n}e^{in\theta},$
the ‘Clebsch-Gordan coefficients’ of $\mathbb{S}_{F}^{1}$ are given by
$c_{mn}^{F~{}~{}p}=\sum_{k,q\in\mathbb{Z}}F^{k}c_{mn}^{~{}~{}q}c_{kq}^{~{}~{}p}=\sum_{k,q\in\mathbb{Z}}F^{k}\delta_{q,m+n}\delta_{p,k+q}=\sum_{k\in\mathbb{Z}}F^{k}\delta_{p,m+n+k}=F^{p-m-n},$
whereas the coefficients $P_{n}^{~{}n^{\prime}}$, generally defined by (2.26), are in this case given by
$P_{n}^{~{}n^{\prime}}=\sum_{m\in\mathbb{Z}}F^{m}c_{mn}^{~{}~{}n^{\prime}}=\sum_{m\in\mathbb{Z}}F^{m}\delta_{n^{\prime},m+n}=F^{n^{\prime}-n},$
such that
$c_{mn}^{F~{}~{}p}=P_{m+n}^{~{}~{}~{}p}.$
. Moreover, the Hilbert basis of $L^{2}\left(\mathbb{S}_{F}^{1}\right)$ (3.10) can be rewritten as
$\mathcal{B}_{F}=\left\{e_{n}^{F}\left(\theta\right)=\sum_{n^{\prime}\in\mathbb{Z}}P_{n}^{~{}n^{\prime}}e_{n^{\prime}}\left(\theta\right),~{}~{}n\in\mathbb{Z}\right\}.$
As in the general treatment of Section 2.3, the natural choice for the Hermitean operator associated to $\mathbb{S}_{F}^{1}$ would be
$d_{F}=-i\frac{1}{\sqrt{F(\theta)}}\partial_{\theta}\sqrt{F(\theta)}\ ,$ (3.11)
but, as treated above, the differential operator that we have to take in duality with the two-cocycle $\omega^{F}$ is the original Hermitean operator $d$, and not the operator $d_{F}$.
By observing that the generators of $\mathfrak{g}\left(\mathbb{S}_{F}^{1}\right)$ are $T_{am}^{F}\equiv\frac{1}{\sqrt{F\left(\theta\right)}}T_{am}$, the two-cocycle of $\mathbb{S}_{F}^{1}$ reads
$\omega_{F}\left(T_{am}^{F},T_{bn}^{F}\right)$ $=$ $\sum_{m^{\prime},n^{\prime}\in\mathbb{Z}}P_{m}^{~{}m^{\prime}}P_{n}^{~{}n^{\prime}}\omega\left(T_{am^{\prime}},T_{bn^{\prime}}\right)=k\eta_{ab}\sum_{m^{\prime},n^{\prime}\in\mathbb{Z}}n^{\prime}\delta_{n^{\prime},-m^{\prime}}F^{m^{\prime}-m}F^{n^{\prime}-n}$ (3.12)
$=$ $k\eta_{ab}\sum_{n^{\prime}\in\mathbb{Z}}n^{\prime}F^{-n^{\prime}-m}F^{n^{\prime}-n}.$
Thus, from Sections 2.2 and 2.3 the corresponding deformation of the affine Lie algebra is given by
$\widetilde{\mathfrak{g}}(\mathbb{S}_{F}^{1})=\Big{\{}\mathcal{T}_{am}^{F},\;d,\;k\Big{\}},$
with Lie brackets reading
$\left[\mathcal{T}_{am}^{F},\mathcal{T}_{bn}^{F}\right]$ $=$ $if_{ab}^{~{}~{}c}\sum_{p\in\mathbb{Z}}c_{mn}^{F~{}~{}p}\mathcal{T}_{cp}^{F}+\omega_{F}\left(\mathcal{T}_{am}^{F},\mathcal{T}_{bn}^{F}\right)=$
$=$ $if_{ab}^{~{}~{}c}\sum_{p\in\mathbb{Z}}F^{p}\mathcal{T}_{c|m+n+p}^{F}+k\eta_{ab}\sum_{n^{\prime}\in\mathbb{Z}}n^{\prime}F^{-n^{\prime}-m}F^{n^{\prime}-n};$
$\left[d,\mathcal{T}_{am}^{F}\right]$ $=$ $\sum_{m^{\prime}\in\mathbb{Z}}P_{m}^{~{}m^{\prime}}m^{\prime}\mathcal{T}_{am^{\prime}}^{F}=\sum_{p\in\mathbb{Z}}\left(m+p\right)F^{p}\mathcal{T}_{a|m+p}^{F}.$ (3.13)
Next, we would like to construct a central extension of the algebra of vector fields on $\mathbb{S}_{F}^{1}$. It is natural to consider $\ell_{m}^{F}(\theta)=\frac{i}{\sqrt{F(\theta)}}e^{im\theta}\partial_{\theta}\sqrt{F(\theta)}$ (see $d_{F}$ in (3.11)). However, in this case the Lie bracket of the vectors $[\ell_{m}^{F},\ell_{n}^{F}]$ on $\mathbb{S}_{F}^{1}$ coincide with the Lie bracket of the de Witt algebra (3.4). This in particular means that the Virasoro algebra associated to the ‘soft’ circle reduces to the usual Virasoro algebra. In other words, we can also see that the Gel’fand–Fuks cocycle $\tilde{\omega}^{F}$ associated to $\{\ell_{n}^{F},n\in\mathbb{\}}$ is given by
$\tilde{\omega}^{F}(e_{n}^{F},e_{m}^{F})=-i\frac{c}{12}\frac{1}{2\pi}\int F(\theta)\text{d}\theta\;e_{n}^{F}(\theta)d_{F}^{3}e_{m}^{F}(\theta)=\omega(e_{n},e_{m})\ .$ (3.14)
We conclude that on the ‘soft’ circle we can define a semi-direct product of the deformed affine Lie algebra (3.13) with the Virasoro algebra (see (3.7) and (3.8)).
The curvature associated to the metric of the ‘soft’ circle is $R=0$. Thus we can perform a global change of variable in (3.14) $\theta\rightarrow\psi$ such that
$F(\theta)=\frac{\text{d}\psi}{\text{d}\theta}\ ,$
and the scalar product (3.14) reduces within the new coordinate $\psi$ to the scalar product on the undeformed circle (3.1).
This in particular means that the semi-direct product of the affine Lie algebra with the Virasoro algebra associated to the ‘soft’ circle is trivially isomorphic to its undeformed analogue. This is not surprising, since the topological classification of one-dimensional closed curves shows that all such curves are topologically equivalent to the circle $\mathbb{S}^{1}$. This equivalence can be formalized by stating that any closed, connected, one-dimensional manifold without boundary is homeomorphic to $\mathbb{S}^{1}$. Specifically, a homeomorphism is a continuous bijection with a continuous inverse, and in the case of one-dimensional closed curves, this means that regardless of how the curve is deformed or embedded in space, it retains the same topological structure as $\mathbb{S}^{1}$. Mathematically, this follows from the fact that the classification of one-dimensional manifolds shows that $\mathbb{S}^{1}$ is the only connected, compact, boundaryless one-dimensional manifold (cf. e.g. [41]). This result is critical in applications across various fields of mathematics and physics. For instance, in knot theory, while different knots (which are embeddings of $\mathbb{S}^{1}$ into three-dimensional space) may not be equivalent in terms of their embeddings, topologically all knots are still homeomorphic to $\mathbb{S}^{1}$. This is because the study of knots focuses on the way $\mathbb{S}^{1}$ is embedded in $\mathbb{R}^{3}$, but the fundamental topological nature of the curve remains the same. Similarly, in string theory, the worldsheet of a closed string is topologically equivalent to $\mathbb{S}^{1}\times\mathbb{R}$, where $\mathbb{S}^{1}$ represents the closed loop of the string at any fixed point in time. The universal property of $\mathbb{S}^{1}$ as the fundamental one-dimensional closed manifold simplifies the analysis of string propagation and interactions (see e.g. [42]). Thus, the circle $\mathbb{S}^{1}$ serves as the canonical model for all one-dimensional closed curves in topology. Of course, this means that there is no need to define analogues of the affine Lie or Virasoro algebras, as we have just seen.
4 ‘Softening’ of $\mathbb{S}^{3}$ and the associated KM algebra
As a second example, we consider now the KM algebra associated to the ‘softening’of the three-sphere $\mathbb{S}^{3}$. We first reproduce the construction of the KM algebra associated to $SU(2)=\mathbb{S}^{3}$ proposed in [20]. As we want to present these results in an explicit way, we reconsider the construction of $\widetilde{\mathfrak{g}}($SU(2))(2))( 2 ) ) in some detail, using a different approach to that adopted in [20], but along the lines of [29].
4.1 The three-sphere $\mathbb{S}^{3}$
The group SU$(2)$ is defined by the set of $2\times 2$ complex matrices
$SU(2)=\Bigg{\{}U=\begin{pmatrix}z_{1}&-\bar{z}_{2}\\ z_{2}&\phantom{-}\bar{z}_{1}\end{pmatrix}:\ \ z_{1},z_{2}\in\mathbb{C}\ ,\ \ |z_{1}|^{2}+|z_{2}|^{2}=1\Bigg{\}}\cong\mathbb{S}^{3}$ (4.1)
The Lie algebra $\mathfrak{su}(2)$ of $SU(2)$ is generated by $J_{0},J_{\pm}$ with Lie brackets:
$\big{[}J_{0},J_{\pm}\big{]}=\pm J_{\pm}\ ,\ \ \big{[}J_{+},J_{-}\big{]}=2J_{0}\ .$
A parameterisation of the three-sphere $\mathbb{S}^{3}$ is given by
$\begin{split}z_{1}=&\cos\frac{\theta}{2}e^{i\frac{\varphi+\psi}{2}}\\ z_{2}=&\sin\frac{\theta}{2}e^{i\frac{\varphi-\psi}{2}}\end{split}$
with [^10] The asymmetry between the $\varphi$ and $\psi$ is resolved by Bargmann [43] (p. 596 eq(4.15)) since he considers $-2\pi\leq\varphi,\psi<2\pi$, but this parameterisation covers twice the three-sphere.
[…]
The left/right invariant vectors fields (obtained by solving (2.7)) are given by
$\begin{array}[]{llllll}L_{\pm}&=&e^{\pm i\psi}\Big{(}-\frac{i}{\sin\theta}\partial_{\varphi}+i\cot\theta\partial_{\psi}\pm\partial_{\theta}\Big{)}\ ,&L_{0}&=&-i\partial_{\psi}\\ R_{\pm}&=&e^{\pm i\varphi}\Big{(}\phantom{-}\;\frac{i}{\sin\theta}\partial_{\psi}-i\cot\theta\partial_{\varphi}\mp\partial_{\theta}\Big{)}\ ,&R_{0}&=&-i\partial_{\varphi}\end{array}$
and satisfy the commutation relations
$\begin{array}[]{llllll}\big{[}L_{0},L_{\pm}\big{]}&=&\pm L_{\pm}\ ,&\big{[}L_{+},L_{-}\big{]}&=&2L_{0}\\[2.0pt] \big{[}R_{0},R_{\pm}\big{]}&=&\pm R_{\pm}\ ,&\big{[}R_{+},R_{-}\big{]}&=&2R_{0}\end{array}$
as well as
$\big{[}L_{a},R_{b}\big{]}=0\ .$
Observe that $R_{\pm}$ and $L_{\pm}$ are related by means of
$e^{\mp i\psi}L_{\pm}+e^{\mp i\varphi}R_{\pm}=0,$
implying that the corresponding left- and right-invariant one-forms will have the same structure, by replacing $\varphi$ by $\psi$ (see Eqs.[4.1] and [4.1]). The Casimir operator takes the form
$Q=-\partial_{\theta}^{2}-\cot\theta\partial_{\theta}-\frac{1}{\sin^{2}\theta}\Big{(}\partial_{\varphi}^{2}+\partial_{\psi}^{2}\Big{)}+2\frac{\cos\theta}{\sin^{2}\theta}\partial_{\varphi}\partial_{\psi}\ .$
The $SU(2)$ right-invariant one-forms (see Eq.[2.8]) read
$\omega_{1}$ $=$ $\sin\psi\text{d}\theta-\cos\psi\sin\theta\text{d}\varphi$
$\omega_{2}$ $=$ $\cos\psi\text{d}\theta+\sin\psi\sin\theta\text{d}\varphi$ (4.2)
$\omega_{3}$ $=$ $\text{d}\psi+\cos\theta\text{d}\varphi,$
and satisfy $d\omega_{i}=\mathbf{-}\epsilon^{jk}{}_{i}\omega_{j}\wedge\omega_{k}$ ($i,j,k=1,2,3$ with summation over repeated indices understood), which are the Maurer-Cartan equations (their right-invariance with respect to $SU(2)$ is discussed in App. A).
The $SU(2)$ right-invariant one-forms (4.1) enable to define the metric on the round three-sphere:
$\text{d}s^{2}$ $=$ $\omega_{1}^{2}+\omega_{2}^{2}+\omega_{3}^{2}$
$=$ $\text{d}\theta^{2}+\text{d}\psi^{2}+\text{d}\varphi^{2}+2\cos\theta\text{d}\psi\text{d}\varphi\ .$
Note also that the one-form $\omega_{3}$ is invariant under $L_{0}$ and that
$\lambda_{1}$ $=$ $\sin\varphi\text{d}\theta-\cos\varphi\sin\theta\text{d}\psi$
$\lambda_{2}$ $=$ $\cos\varphi\text{d}\theta+\sin\varphi\sin\theta\text{d}\psi$ (4.3)
$\lambda_{3}$ $=$ $\cos\theta\text{d}\psi+\text{d}\varphi$
are left-invariant one-forms, and $\lambda_{3}$ is invariant under $R_{0}$. Of course, we have
$\text{d}s^{2}=\lambda_{1}^{2}+\lambda_{2}^{2}+\lambda_{3}^{2}=\omega_{1}^{2}+\omega_{2}^{2}+\omega_{3}^{2}\ ,$
and thus the metric is left and right invariant, as expected by compactness, reproducing the results above.
Introducing the metric tensor
$g=\begin{pmatrix}1&0&0\\ 0&1&\cos\theta\\ 0&\cos\theta&1\end{pmatrix}$
we define the (normalized) invariant measure of integration on the three-sphere:
$\int\limits_{\mathbb{S}^{3}}\text{d}\mu(\mathrm{SU}(2))=\frac{1}{16\pi^{2}}\int_{\mathbb{S}^{3}}\sqrt{\det g}\;\text{d}\theta\;\text{d}\varphi\;\text{d}\psi=\frac{1}{16\pi^{2}}\int\limits_{0}^{\pi}\sin\theta\;\text{d}\theta\int\limits_{0}^{2\pi}\text{d}\varphi\int\limits_{0}^{4\pi}\text{d}\psi\ .$ (4.4)
4.1.1 Matrix elements
Recall that the unitary representations of SU$(2)$ are given by $\mathcal{D}_{s}=\big{\{}\left|s,n\right>,-s\leq n\leq s\big{\}},s\in\frac{1}{2}\mathbb{N}$ and that we have
$J_{\pm}\big{|}s,n\big{>}$ $=$ $\sqrt{(s\mp n)(s\pm n+1)}\big{|}s,n\pm 1\big{>}$
$J_{0}\big{|}s,n\big{>}$ $=$ $n\big{|}s,n\big{>}$ (4.5)
$Q\big{|}s,n\big{>}$ $=$ $s(s+1)\big{|}s,n\big{>}$
To compute the normalized matrix elements $\psi_{nsm},-s\leq n,m\leq s$ of the representation $\mathcal{D}_{s}$ we proceed as in [29].
  1. 1.
    We solve the differential equations
    $L_{0}\psi_{nsm}(\theta,\varphi,\psi)$ $=$ $n\psi_{nsm}(\theta,\varphi,\psi)$
    $R_{0}\psi_{nsm}(\theta,\varphi,\psi)$ $=$ $m\psi_{nsm}(\theta,\varphi,\psi)$
    $Q\psi_{nsm}(\theta,\varphi,\psi)$ $=$ $s(s+1)\psi_{nsm}(\theta,\varphi,\psi)\ .$
    The first two equations lead obviously to
    $\psi_{nsm}(\theta,\varphi,\psi)$ $=$ $e^{in\psi+im\varphi}F_{nsm}(\theta)\ .$
    The last equation is solved using an appropriate ansatz which depends on the value of $m,n$ (see below and [29]) and expresses the matrix elements in terms of hypergeometric polynomials. At this point the matrix elements are defined up to a constant $C_{nsm}$.
  2. 2.
    We impose that the matrix elements satisfy both for the left and the right action of the Lie algebra $\mathfrak{su}(2)$ relations (4.1.1). This fixes the constant $C_{nsm}$ up to a global factor $C_{s}$.
  3. 3.
    Using the integration (4.4) we impose the normalization condition
    $\big{\|}\psi_{nsm}\big{\|}^{2}=\int\limits_{\mathbb{S}^{3}}\text{d}\mu(\mathrm{SU}(2))\;\big{|}\psi_{nsm}(\theta,\varphi,\psi)\big{|}^{2}=1\ .$
    We recall that we have for $a,b\in\mathbb{N}$
    $\int\limits_{0}^{\pi}\text{d}\theta\;\sin\theta\cos^{2a}\frac{\theta}{2}\sin^{2b}\frac{\theta}{2}=2\frac{a!b!}{(a+b+1)!}\ .$
We thus obtain
$\psi_{nsm}(\theta,\varphi,\psi)=\left\{\begin{array}[]{lc}\frac{(-1)^{m-n}\sqrt{(2s+1)}}{(n-m)!}\sqrt{\frac{(s+n)!}{(s-n)!}\frac{(s-m)!}{(s+m)!}}e^{im\varphi+in\psi}\cos^{-n-m}\frac{\theta}{2}\sin^{n-m}\frac{\theta}{2}&n\geq m\\ {}_{2}F_{1}(-m-s,-m+s+1;1-m+n;\;\sin^{2}\frac{\theta}{2})&-n-m\geq 0\\[6.0pt] \frac{\sqrt{(2s+1)}}{(m-n)!}\sqrt{\frac{(s+m)!}{(s-m)!}\frac{(s-n)!}{(s+n)!}}e^{im\varphi+in\psi}\cos^{-n-m}\frac{\theta}{2}\sin^{m-n}\frac{\theta}{2}&m\geq n\\ {}_{2}F_{1}(-n-s,-n+s+1;1-n+m;\;\sin^{2}\frac{\theta}{2})&-n-m\geq 0\\[6.0pt] \frac{(-1)^{m-n}\sqrt{2s+1}}{(n-m)!}\sqrt{\frac{(s+n)!}{(s-n)!}\frac{(s-m)!}{(s+m)!}}e^{im\varphi+in\psi}\cos^{n+m}\frac{\theta}{2}\sin^{n-m}\frac{\theta}{2}&n\geq m\\ {}_{2}F_{1}(n-s,n+s+1;1+n-m;\;\sin^{2}\frac{\theta}{2})&n+m\geq 0\\[6.0pt] \frac{\sqrt{2s+1}}{(m-n)!}\sqrt{\frac{(s+m)!}{(s-m)!}\frac{(s-n)!}{(s+n)!}}e^{im\varphi+in\psi}\cos^{n+m}\frac{\theta}{2}\sin^{m-n}\frac{\theta}{2}&m\geq n\\ {}_{2}F_{1}(m-s,m+s+1;1+m-n;\;\sin^{2}\frac{\theta}{2})&n+m\geq 0\end{array}\right.$ (4.14)
where ${}_{2}F_{1}$ denotes the Euler hypergeometric polynomial (see e.g. [29] for definition). Under these conditions, the set
=$\mathcal{B}=\Big{\{}\psi_{nsm},s\in\frac{1}{2}\mathbb{N},-s\leq n,m\leq s\Big{\}}$ (4.15)
constitutes a Hilbert basis of $L^{2}(\mathrm{SU}(2))$ and we have:
$(\psi_{nsm},\psi_{n^{\prime},s^{\prime},m^{\prime}})=\int\limits_{\mathbb{S}^{3}}\text{d}\mu(\mathrm{SU}(2))\;\overline{\psi_{nsm}}(\theta,\varphi,\psi)\psi_{n^{\prime}s^{\prime}m^{\prime}}(\theta,\varphi,\psi)=\delta_{ss^{\prime}}\delta_{nn^{\prime}}\delta_{mm^{\prime}}\ .$
4.1.2 Clebsch-Gordan coefficients
Considering the tensor product
$\mathcal{D}_{s_{1}}\otimes\mathcal{D}_{s_{2}}=\bigoplus\limits_{S=|s_{1}-s_{2}|}^{s_{1}+s_{2}}\mathcal{D}_{S}\ ,$
and introducing the Clebsch-Gordan coefficients $\begin{pmatrix}s_{1}&s_{2}&S\\ m_{1}&m_{2}&m_{1}+m_{2}\end{pmatrix}$ we have
$\big{|}S,m_{1}+m_{2}\big{>}=\sum\limits_{S=|s_{1}-s_{2}|}^{s_{1}+s_{2}}{\scriptsize\begin{pmatrix}s_{1}&s_{2}&S\\ m_{1}&m_{2}&m_{1}+m_{2}\end{pmatrix}\mid s_{1},m_{1}\rangle\otimes\mid s_{2},m_{2}\rangle\ .}$
Since our matrix elements are normalized such that
$\psi_{nsm}(0,0,0)=\sqrt{2s+1}$
we obtain
$\psi_{n_{1}s_{1}m_{1}}(\theta,\varphi,\psi)\psi_{n_{2}s_{2}m_{2}}(\theta,\varphi,\psi)=\sum\limits_{S=|s_{1}-s_{2}|}^{s_{1}+s_{2}}c_{s_{1}s_{2}}^{S}{}_{n_{1}n_{2}m_{1}m_{2}}\;\psi_{n_{1}+n_{2}Sm_{1}+m_{2}}(\theta,\varphi,\psi)$
with
$c_{s_{1}s_{2}}^{S}{}_{n_{1}n_{2}m_{1}m_{2}}\equiv\sqrt{\frac{(2s_{1}+1)(2s_{2}+1)}{2S+1}}\begin{pmatrix}s_{1}&s_{2}&S\\ n_{1}&n_{2}&n_{1}+n_{2}\end{pmatrix}\;\overline{\begin{pmatrix}s_{1}&s_{2}&S\\ m_{1}&m_{2}&m_{1}+m_{2}\end{pmatrix}}$ (4.16)
4.2 A KM algebra associated to $\mathbb{S}^{3}$
The KM algebra associated to SU$(2)=\mathbb{S}^{3}$ follows directly from (4.16) and [20]. Indeed we have
$\widetilde{\mathfrak{g}}(\text{SU}(2))=\Big{\{}\mathcal{T}_{ansm},L_{0},R_{0},k_{L},k_{R},\ \ a=1,\cdots,d\ ,\ell\in\frac{1}{2}\mathbb{N},-\ell\leq n,m\leq\ell\Big{\}}$
with the central charges $k_{L},k_{R}$ associated to the two-cocycles:
$\omega_{L}(x,y)$ $=$ $k_{L}\int\limits_{\mathbb{S}^{3}}\text{d}\mu(\mathrm{SU}(2))\;\big{<}x,L_{0}y\big{>}_{0}$
$\omega_{R}(x,y)$ $=$ $k_{R}\int\limits_{\mathbb{S}^{3}}\text{d}\mu(\mathrm{SU}(2))\;\big{<}x,R_{0}y\big{>}_{0}$
The Lie brackets take then the form (see [20])
$\big{[}\mathcal{T}_{ansm},\mathcal{T}_{a^{\prime}n^{\prime}s^{\prime}m^{\prime}}\big{]}$ $=$ $if_{aa^{\prime}}{}^{a^{\prime\prime}}c_{ss^{\prime}nn^{\prime}mm^{\prime}}^{s^{\prime\prime}}\mathcal{T}_{a^{\prime\prime}n+n^{\prime}s^{\prime\prime}m+m^{\prime}}$
$+\eta_{ab}(-1)^{m-n}\delta_{ss^{\prime}}\delta_{n,-n^{\prime}}\delta_{m,-m^{\prime}}(k_{L}n^{\prime}+k_{R}m^{\prime})$
$\big{[}L_{0},\mathcal{T}_{ansm}\big{]}$ $=$ $n\mathcal{T}_{ansm}$
$\big{[}R_{0},\mathcal{T}_{ansm}\big{]}$ $=$ $m\mathcal{T}_{ansm}$
4.3 A KM algebra associated to $\mathbb{S}^{3}_{F}$
Consider now a ‘softened’ three-sphere $\mathbb{S}_{F}^{3}$ with a deformed metric $g_{MN}^{\mu}$ such that $\sqrt{\det{(g_{MN}^{\mu})}}=F(\theta,\psi,\varphi)\sin\theta$. The scalar product on $L^{2}(\mathbb{S}_{F}^{3})$ reduces to
$(f,g)_{F}=\frac{1}{16\pi^{2}}\int\limits_{0}^{\pi}\text{d}\theta\int\limits_{0}^{2\pi}\text{d}\varphi\int\limits_{0}^{4\pi}\text{d}\psi\;\sin\theta\;F(\theta,\psi,\varphi)\;\overline{f(\theta,\psi,\varphi)}g(\theta,\psi,\varphi)\ .$
The results of Sections 2.2 and 2.3 enable us to define KM algebras associated to $\mathbb{S}_{F}^{3}$:
  1. 1.
    The Hilbert basis of $L^{2}(\mathbb{S}_{F}^{3})$ is
    $\mathcal{B}_{F}=\Big{\{}\psi_{nsm}^{F}(\theta,\psi,\varphi)=\frac{1}{\sqrt{F\left(\theta,\psi,\varphi\right)}}\psi_{nsm}\left(\theta,\psi,\varphi\right)=P_{nsm}^{~{}~{}~{}prq}\psi_{prq}(\theta,\psi,\varphi)\ ,\ \ s\in\frac{1}{2}\mathbb{N},-s\leq n,m\leq s\Big{\}},$
    where the coefficients $P_{nsm}^{~{}~{}~{}prq}$ are defined according to the general formula (2.26).
  2. 2.
    The Hermitean operators associated to $\mathbb{S}_{F}^{3}$ are $L_{0}$ and $R_{0}$ themselves.
  3. 3.
    The two-cocycles associated to $L_{0}$ and $R_{0}$ are given by
    $\omega_{\alpha}^{F}\left(T_{ansm}^{F},T_{bn^{\prime}s^{\prime}m^{\prime}}^{F}\right)$ $=$ $P_{nsm}^{~{}~{}~{}prq}P_{n^{\prime}s^{\prime}m^{\prime}}^{~{}~{}~{}~{}~{}p^{\prime}r^{\prime}q^{\prime}}\omega_{\alpha}\left(T_{aprq},T_{bp^{\prime}r^{\prime}q^{\prime}}\right)$
    $=$ $\eta_{ab}\left(-1\right)^{m-n}P_{nsm}^{~{}~{}~{}prq}P_{n^{\prime}s^{\prime}m^{\prime}}^{~{}~{}~{}~{}~{}p^{\prime}r^{\prime}q^{\prime}}\delta_{ss^{\prime}}\delta_{n,-n^{\prime}}\delta_{m,-m^{\prime}}\left(\delta_{\alpha L}k_{L}p^{\prime}+\delta_{\alpha R}k_{R}q^{\prime}\right),$
    where $\alpha=L$ and $R$, respectively.
Thus, from Sections 2.2 and 2.3 the KM algebra associated to the ‘softened’ three-sphere $\mathbb{S}_{F}^{3}$ is
$\widetilde{\mathfrak{g}}(\mathbb{S}_{F}^{3})=\Big{\{}\mathcal{T}_{ansm}^{F},L_{0},R_{0},k_{L},k_{R}\ ,\ \ a=1,\cdots,d,s\in\frac{1}{2}\mathbb{N},-s\leq n,m\leq s\Big{\}}$
Taking into account that
$\frac{1}{\sqrt{F(\theta,\psi,\varphi)}}=\sum_{s\in\frac{1}{2}\mathbb{N}}\sum\limits_{n=-s}^{s}\sum\limits_{m=-s}^{s}F^{nsm}\psi_{nsm}(\theta,\psi,\varphi),$
the Lie brackets reduce to
$\big{[}\mathcal{T}_{ansm}^{F},\mathcal{T}_{a^{\prime}n^{\prime}s^{\prime}m^{\prime}}^{F}\big{]}$ $=$ $if_{aa^{\prime}}{}^{a^{\prime\prime}}c_{ss^{\prime}nn^{\prime}mm^{\prime}}^{Fs^{\prime\prime}}\mathcal{T}_{a^{\prime\prime}n+n^{\prime}s^{\prime\prime}m+m^{\prime}}^{F}$ (4.17)
$+\eta_{ab}\left(-1\right)^{p-q}P_{nsm}^{~{}~{}~{}prq}P_{n^{\prime}s^{\prime}m^{\prime}}^{~{}~{}~{}p^{\prime}r^{\prime}q^{\prime}}\delta_{rr^{\prime}}\delta_{p,-p^{\prime}}\delta_{q,-q^{\prime}}\left(k_{L}p^{\prime}+k_{R}q^{\prime}\right);$
$\big{[}L_{0},\mathcal{T}_{ansm}^{F}\big{]}$ $=$ $pP_{nsm}^{~{}~{}~{}prq}\mathcal{T}_{aprq}^{F};$
$\big{[}R_{0},\mathcal{T}_{ansm}^{F}\big{]}$ $=$ $qP_{nsm}^{~{}~{}~{}prq}\mathcal{T}_{aprq}^{F}.$
where the ‘Clebsch-Gordan coefficients’ of $\mathbb{S}_{F}^{3}$ $c_{ss^{\prime}nn^{\prime}mm^{\prime}}^{Fs^{\prime\prime}}$ are defined by an expression analogous to (2.25).
4.3.1 $\widetilde{\mathbb{S}}^{3}$ and its physical applications
An important example of ‘softened’ three-sphere is the Berger three-sphere, also named squashed three-sphere, denoted by $\widetilde{\mathbb{S}}^{3}$. Introducing the right-invariant one-forms (4.1) and following Berger [44], the (doubly) squashed three-sphere $\widetilde{\mathbb{S}}^{3}$ is endowed with the following metric:
$\text{d}s^{2}$ $=$ $\omega_{1}^{2}+b^{2}\omega_{2}^{2}+c^{2}\omega_{3}^{2}$ (4.18)
$=$ $\Big{(}\sin^{2}\psi+b^{2}\cos^{2}\psi\Big{)}\text{d}\theta^{2}+c^{2}\text{d}\psi^{2}+\Big{(}\sin^{2}\theta\cos^{2}\psi+b^{2}\sin^{2}\psi\sin^{2}\theta+c^{2}\cos^{2}\theta\Big{)}\text{d}\varphi^{2}$
$+2(b^{2}-1)\sin\psi\cos\psi\sin\theta\text{d}\theta\text{d}\varphi+2c^{2}\cos^{2}\theta\text{d}\varphi\text{d}\psi\ ,$
where $b,c>0$ are named squashing parameters. Thus, the metric tensor takes the form
$g=\begin{pmatrix}\sin^{2}\psi+b^{2}\cos^{2}\psi&0&(1-b^{2})\sin\psi\cos\psi\sin\theta\\ 0&{c}^{2}&c^{2}\cos\theta\\ (b^{2}-1-1)\sin\psi\cos\psi\sin\theta&{c}^{2}\cos\theta&\sin^{2}\theta\cos^{2}\psi+b^{2}\sin^{2}\psi\sin^{2}\theta+c^{2}\cos^{2}\theta\end{pmatrix}$
and
$\sqrt{\det g}=bc\sin\theta\ .$
As $\omega_{1},\omega_{2},\omega_{3}$ in (4.1) are right-invariant (but not left-invariant) one-forms, the metric (4.18) is invariant under the right-action of SU$(2)$, but not under the left-action of SU$(2)$. Thus, $R_{\pm},R_{0}$ generate isometries of the squashed sphere. Moreover, the scalar curvature of $\widetilde{\mathbb{S}}^{3}$ is given by
$R=-\frac{b^{4}+(c^{2}-1)^{2}-2b^{2}(c^{2}+1)}{2b^{2}c^{2}},$
which for suitable choices of $b$ and $c$, can be made negative; for instance, by setting $b=1$, one obtains
$R=2-c^{2}/2<0\Leftrightarrow c^{2}>4\ .$
The squashed three-sphere $\widetilde{\mathbb{S}}^{3}$ with two squashing parameters $b$ and $c$ provides a rich framework for understanding various physical phenomena. Indeed, $b$ and $c$ introduce anisotropic scaling in the directions corresponding to the $SU(2)$ right-invariant one-forms $\omega_{2}$ and $\omega_{3}$, and therefore the original $SO(4)$ isometry of $\mathbb{S}^{3}$ gets reduced, e.g. typically to $SU(2)$, for generic values of $b$ and $c$.
In the mathematics context, $\widetilde{\mathbb{S}}^{3}$ was considered by Hitchin [45] in his discussion of the space of harmonic spinors (i.e., the null space of the Dirac operator) on a manifold: in fact, $\widetilde{\mathbb{S}}^{3}$ is a notable illustration of the fact that the number of harmonic spinors is not a topological invariant of the manifold, but rather it depends on the particular metric, as well. In physics, scalar quantum field theory on $\widetilde{\mathbb{S}}^{3}$ has been investigated by several authors [46, 47, 48], mainly due to its appearance as (a particular case of) the spatial section of the mixmaster cosmological model [49].
Moreover, $\widetilde{\mathbb{S}}^{3}$ naturally arises in flux compactifications and Kaluza-Klein (KK) reductions. For example, compactifications of 11-dimensional supergravity or 10-dimensional string theory on $\widetilde{\mathbb{S}}^{3}$ lead to modifications in the low-energy effective theory, where the parameters $b$ and $c$ break the internal symmetry, affecting the spectrum of KK modes [50, 51, 19]. As first discussed in [52], these deformations can break part of the supersymmetry and affect the vacuum structure, the cosmological constant and the gauge couplings in the lower-dimensional theory. Within the AdS/CFT correspondence, $\widetilde{\mathbb{S}}^{3}$ plays a significant role, for instance in the AdS3 solutions of supergravity and string theory, since the geometry of the internal manifold, controlled by the squashing parameters, affects the dual two-dimensional conformal field theory (CFT). On the other hand, within the gauge/gravity duality, the gravity duals of supersymmetric gauge theories on $\widetilde{\mathbb{S}}^{3}$ (with various types of squashing) have been investigated [53, 54], as well. Compactifications involving $\widetilde{\mathbb{S}}^{3}$ are also important in M-theory, because $\widetilde{\mathbb{S}}^{3}$ can form part of the internal spaces with $\mathit{G}_{2}$ holonomy, as seen in supergravity solutions with fluxes. It is also worth recalling here that the effective actions for both scalars and fermions on $\widetilde{\mathbb{S}}^{3}$ have been obtained in [55, 56], with the aim to compare with the AdS/CFT results of [57, 58, 59]; notice that the relation is far from being obvious, since the regimes where the results are expected to apply are very different (cf. e.g. [60] for a discussion).
We should recall that $\widetilde{\mathbb{S}}^{3}$ is also a key tool in localization techniques within the aforementioned supersymmetric gauge theories; specifically, in three-dimensional $\mathcal{N}=2$ supersymmetric gauge theories, placing the theory on $\widetilde{\mathbb{S}}^{3}$ allows for the exact computation of partition functions and other observables through localization. Hama, Hosomichi, and Lee [61] showed that the squashing parameters $b$ and $c$ modify the background geometry, influencing the preserved supersymmetry. This affects the localization locus of the path integral, which in turn modifies the resulting physical observables, such as the exact partition function and the Wilson loops. These computations are particularly useful for studying non-perturbative effects and for testing dualities, such as mirror symmetry and (three-dimensional) Seiberg-like dualities. In general, $b$ and $c$ induce a continuous deformation of the background geometry, thus enabling the investigation of different phases of the theory. For instance, in supersymmetric gauge theories with exact localization, the partition function is an integral over the moduli space of flat connections, and the squashing parameters alter the effective action and measure of the integral, as noted again in [61].
The invariant measure on the squashed sphere reads
$\int\limits_{{\mathbb{S}^{3}}}\text{d}\mu(\widetilde{\mathbb{S}^{3}})=\frac{1}{16\pi^{2}}\int\limits_{{\mathbb{S}^{3}}}\sqrt{\det g}\;\text{d}\theta\;\text{d}\varphi\;\text{d}\psi=\frac{1}{16\pi^{2}}\int\limits_{0}^{\pi}cb\sin\theta\;\text{d}\theta\int\limits_{0}^{2\pi}\text{d}\varphi\int\limits_{0}^{4\pi}\text{d}\psi\ .$ (4.19)
This means that
$\mathcal{B}_{\widetilde{\mathbb{S}}^{3}}=\Big{\{}\widetilde{\psi}_{nsm}=\frac{1}{\sqrt{bc}}\psi_{nsm}\ ,\ \ s\in\frac{1}{2}\mathbb{N}\ ,\ \ -s\leq n,m\leq s\Big{\}}$
with $\psi_{nsm}$ corresponding to the matrix elements of SU$(2)$ (see e.g. (4.15)) constitutes a Hilbert basis of the squashed sphere:
$(\widetilde{\psi}_{nsm},\widetilde{\psi}_{n^{\prime}s^{\prime}m^{\prime}})=\int\limits_{{\mathbb{S}^{3}}}\text{d}\mu(\widetilde{\mathbb{S}^{3}})\;\overline{\widetilde{\psi}_{nsm}}(\theta,\varphi,\psi)\;\widetilde{\psi}_{n^{\prime}s^{\prime}m^{\prime}}(\theta,\varphi,\psi)=\delta_{ss^{\prime}}\delta_{nn^{\prime}}\delta_{mm^{\prime}}\ .$
Note that, as mentioned in Section 4.2, $\widetilde{\psi}_{nsm}$ are in representation of the left and right actions of the group SU$(2)$, but only the right action is an isometry of the squashed sphere.
From Section 4.1.2 the product $\widetilde{\psi}_{nsm}(\theta,\psi,\phi)\widetilde{\psi}_{n^{\prime}s^{\prime}m^{\prime}}(\theta,\psi,\phi)$ is straightforward. The KM algebra associated to the squashed sphere $\widetilde{\mathfrak{g}}(\widetilde{\mathbb{S}}^{3})$ follows directly, and it is isomorphic to the KM algebra associated to the usual three-sphere. Since the usual sphere is obtained taking the limit $b,c\to 1$, considering this limit we recover an isomorphic realization of the KM algebra of $\mathbb{S}^{3}$ from the KM algebra of the squashed sphere.
5 ‘Softening’ of $\mathbb{S}^{2}$ and the associated KM algebra
As a third example, we consider now the KM algebra associated to the ‘softening’of the two-sphere $\mathbb{S}^{2}$.
5.1 The two-sphere $\mathbb{S}^{2}$
The two-sphere $\mathbb{S}^{2}$ is given by the symmetric space $SO(3)/SO(2)\simeq SU(2)/U(1)=\mathbb{C}P^{1}$, namely by the complex projective line (here $\simeq$ denotes isomorphism of homogeneous spaces). The complex projective line $\mathbb{C}P^{1}$ has profound applications in several areas of theoretical physics, particularly in gauge theory, string theory, and twistor theory, because the isomorphism $\mathbb{C}P^{1}\simeq\mathbb{S}^{2}$ makes it a useful model for understanding internal symmetries and topological properties in physical systems. In gauge theory, for instance, $\mathbb{C}P^{1}$ plays a crucial role in the study of monopole solutions. The celebrated ’t Hooft-Polyakov monopole solution, which arises in non-Abelian gauge theory, uses $\mathbb{C}P^{1}$ as the internal symmetry space that describes the direction of the Higgs field at spatial infinity. The mapping from spatial infinity $\mathbb{S}^{2}$ to $\mathbb{C}P^{1}$, classified by the homotopy group $\pi_{2}(SU(2)/U(1))\cong\mathbb{Z}$, gives rise to the topological charge of the monopole, which corresponds to the magnetic charge [62, 63].
In string theory, $\mathbb{CP}^{1}$ arises naturally in sigma models, where it serves as the target space of two-dimensional field theories describing strings. On the other hand, in topological string theory $\mathbb{C}P^{1}$ provides a simple setting for computing Gromov-Witten invariants, counting the number of holomorphic maps from the string worldsheet to the target space (see e.g. [64]). Additionally, $\mathbb{C}P^{1}$ often appears as part of the internal geometry in compactifications of higher-dimensional theories, such as compactifications of type II string theory on Calabi-Yau manifolds, where $\mathbb{CP}^{1}$ can represent two-cycles within the compactified space, in turn determining the low-energy effective theory and the spectrum of BPS states.
Furthermore, in twistor theory $\mathbb{C}P^{1}$ appears in the context of describing spacetime in terms of complex geometry. Indeed, the twistor space $\mathbb{C}P^{3}$, which encodes information about the four-dimensional Minkowski spacetime, is fibered over $\mathbb{CP}^{1}$, with each point in spacetime corresponding to a projective line in twistor space. This reformulation of spacetime, where points are replaced by lines in a complex projective space, allows the description of gravitational and gauge field theories in terms of holomorphic structures. The incidence relation between points in twistor space and lines in $\mathbb{C}P^{1}$ is given by a simple geometric condition, leading to the so-called Penrose transform, converting solutions of wave equations in spacetime into holomorphic functions in twistor space [65]. This geometric framework has been successfully applied to study the scattering amplitudes of gauge theories and gravity, providing new insights into the structure of field theories.
In the following, we assume that $Q\in U(1)\subset SU(2)$ is given by
$Q=e^{i\theta R_{0}}\ .$
The points $(\theta,\psi,\varphi=\text{cons.})$. ) parameterise points on the manifold $\mathbb{S}^{2}\cong SU(2)/U(1)\subset\mathbb{S}^{3}\cong SU(2).$ With this parameterisation, for the level surfaces $\varphi=$ cons. we have, on the one hand
$L_{\pm}=e^{\pm i\psi}\Big{(}i\cot\theta\partial_{\psi}\pm\partial_{\theta}\Big{)}\ ,\ \ L_{0}=-i\partial_{\psi}\ ,$
as well as the relation
$(f,g)$ $=$ $\frac{1}{16\pi^{2}}\int\limits_{0}^{\pi}\sin\theta\text{d}\theta\int\limits_{0}^{2\pi}\text{d}\varphi\int\limits_{0}^{4\pi}\text{d}\psi\;\overline{f(\theta,\psi,\varphi)}\;{g(\theta,\psi,\varphi)}$
$=$ $\Big{|}_{\varphi=\text{cons.}}\;\frac{1}{8\pi}\int\limits_{0}^{\pi}\sin\theta\;\text{d}\theta\int\limits_{0}^{4\pi}\text{d}\psi\overline{f\theta,\psi,\varphi=\text{cons.})}\;{g(\theta,\psi,\varphi=\text{cons.})}\ .$. ) .
The overall factor is due to the fact that we cover twice the two-sphere. From now on we set $\psi\in[0,2\pi[$ and substitute the normalization factor in the integral above by $1/4\pi$. Since the harmonic analysis on the SU(2)/(2)/( 2 ) /U$(1)$ coset (and the related matrix elements) must be U$(1)-$invariant, the only matrix elements which appear are those which are $U(1)-$invariant, thus corresponding to the matrix elements with $m=0$ (bosonic case with $s\in\mathbb{N}$), which in turn reduce to spherical harmonics. We thus have the explicit polynomial expression for spherical harmonics (the sign $(-1)^{s}$ below is to agree with standard definition of spherical harmonics):
$\begin{array}[]{c}Y_{sn}(\theta,\psi)=\frac{(-1)^{{s}+\frac{|n|+n}{2}}\sqrt{2s+1}}{|n|!}\sqrt{\frac{(s+|n|)!}{(s-|n|)!}}e^{in\psi}\cos^{|n|}\frac{\theta}{2}\sin^{|n|}\frac{\theta}{2}{}_{2}F_{1}(|n|-s,|n|+s+1;1+|n|;\;\sin^{2}\frac{\theta}{2})\end{array}$ (5.2)
Thus $\{Y_{sn},s\in\mathbb{N},-s\leq n\leq s\}$ constitutes a Hilbert basis of $L^{2}(\mathbb{S}^{2})$ and we have
$(Y_{sn},Y_{s^{\prime}n^{\prime}})=\delta_{ss^{\prime}}\delta_{nn^{\prime}}\ .$
As in Section 4.1.2, we also have
$Y_{s_{1}n_{1}}(\theta,\psi)Y_{s_{2}n_{2}}(\theta,\psi)=\sum\limits_{S=|s_{1}-s_{2}|}^{s_{1}+s_{2}}c_{s_{1}s_{2}}^{S}{}_{n_{1}n_{2}}\;Y_{Sn_{1}+n_{2}}(\theta,\psi)$ (5.3)
with
$c_{s_{1}s_{2}}^{S}{}_{n_{1}n_{2}}\equiv\sqrt{\frac{(2s_{1}+1)(2s_{2}+1)}{2S+1}}\begin{pmatrix}s_{1}&s_{2}&S\\ n_{1}&n_{2}&n_{1}+n_{2}\end{pmatrix}\;\begin{pmatrix}s_{1}&s_{2}&S\\ 0&0&0\end{pmatrix}$ (5.4)
5.2 KM and Virasoro algebras associated to $\mathbb{S}^{2}$
The KM algebra associated to $\mathbb{S}^{2}$ follows directly from (5.4) and the results in [20]. Indeed, we have
$\widetilde{\mathfrak{g}}(\mathbb{S}^{2})=\Big{\{}\mathcal{T}_{a\ell m},J_{0},k\ ,a=1,\cdots\dim\mathfrak{g}\ ,\ell\in\mathbb{N}\ ,-\ell\leq m\leq\ell\Big{\}}\ ,$
with the central charge $k$ associated to the two-cocycle:
$\omega(x,y)=\frac{k}{4\pi}\int\limits_{0}^{\pi}\sin\theta\;\text{d}\theta\int\limits_{0}^{2\pi}\text{d}\psi\;\big{<}x,J_{0}y\big{>}_{0}\ .$
The Lie brackets take the form (see [20])
$\big{[}\mathcal{T}_{a_{1}\ell_{1}m_{1}},\mathcal{T}_{a_{2}\ell_{2}m_{2}}\big{]}$ $=$ $if_{a_{1}a_{2}}{}^{a_{3}}c_{\ell_{1},m_{1},\ell_{2},m_{2}}^{\ell_{3}}\mathcal{T}_{a_{3}\ell_{3}m_{1}+m_{2}}+(-1)^{m_{1}}km_{2}\;\eta_{a_{1}a_{2}}\delta_{\ell_{1},\ell_{2}}\delta_{m_{1},-m_{2}}\ ,$ (5.5)
$\big{[}J_{0},\mathcal{T}_{a\ell m}\big{]}$ $=$ $m\mathcal{T}_{a\ell m}\ .$
An analogue of the Virasoro algebra on the two-sphere was defined in [24]. We briefly reproduce here the results in a way suitable to extend this algebra on the ‘softened’ sphere. To proceed, set
$Y_{\ell m}(\theta,\psi)=Q_{\ell m}(\theta)e^{im\psi}\ .$
Note also that for $f\in L^{2}(\mathbb{S}^{2})$ we have
$f(\theta,\psi)=\sum\limits_{\ell=0}^{+\infty}\sum\limits_{m=-\ell}^{\ell}f^{\ell m}Y_{\ell m}(\theta,\psi)=\sum\limits_{m\in\mathbb{Z}}\Big{(}\sum\limits_{\ell\geq|m|}f^{\ell m}Q_{\ell m}(\theta)\Big{)}e^{im\psi}\ .$
The second summation is not usual, but more appropriate on purpose. The orthogonality between spherical harmonics implies
$\frac{1}{2}\int\limits_{0}^{2\pi}\sin\theta\;\text{d}\theta\;Q_{\ell m}(\theta)Q_{\ell^{\prime}m}(\theta)=\delta_{\ell\ell^{\prime}}$
and (5.3) leads to
$Q_{\ell m}(\theta)Q_{\ell^{\prime}m^{\prime}}(\theta)=c_{\ell,m,\ell^{\prime},m^{\prime}}^{\ell^{\prime\prime}}Q_{\ell^{\prime\prime},m+m^{\prime}}(\theta)\ .$
We now introduce
$\ell_{\ell m}=iQ_{\ell m}(\theta)e^{im\psi}\partial_{\psi}\ .$ (5.6)
The set $\{\ell_{\ell m},m\in\mathbb{N},\ell\geq|m|\}$ constitutes a subset of the vector fields on the two-sphere with Lie brackets:
$\big{[}\ell_{\ell m},\ell_{\ell^{\prime}m^{\prime}}\big{]}=(m-m^{\prime})c_{\ell,m,\ell^{\prime},m^{\prime}}^{\ell^{\prime\prime}}\ell_{\ell^{\prime},m+m^{\prime}}\ .$ (5.7)
This algebra admits a non-trivial two-cocycle, which is analogous to the Gel’fand–Fuks cocycle for the Virasoro algebra:
$\omega(f,g)=-\frac{i}{12}\frac{c}{4\pi}\int\limits_{0}^{\pi}\sin\theta\text{d}\theta\int\limits_{0}^{2\pi}\text{d}\psi f(\theta,\psi)\partial_{\psi}^{3}g(\theta,\psi)\ .$
Evaluated on the vector fields (5.7), this gives
$\omega(\ell_{\ell,m},\ell_{\ell^{\prime},m^{\prime}})=\frac{c}{12}m^{3}\delta_{\ell\;\ell^{\prime}}\delta_{m,-m^{\prime}}\ .$ (5.8)
Let $\mathcal{L}_{\ell m}$ be the generators of the centrally extended algebra (5.7) by the cocycle (5.8). Define now $L_{\ell m}=\mathcal{L}_{\ell m}+\frac{c}{24}\delta_{m,0}$. The Virasoro algebra of the two-sphere
$\mathrm{Vir}(\mathbb{S}^{2})=\Big{\{}L_{\ell,m}\ ,\ \ m\in\mathbb{Z}\ ,\ \ \ell\geq|m|\ ,\ \ c\Bigg{\}}$
has Lie brackets
$\big{[}L_{\ell m},L_{\ell^{\prime}m^{\prime}}\big{]}=(m-m^{\prime})\;c_{\ell,m,\ell^{\prime},m^{\prime}}^{\ell^{\prime\prime}}L_{\ell^{\prime\prime}m+m^{\prime}}+\frac{c}{12}(m^{3}-m)\delta_{\ell\ell^{\prime}}\delta_{m,-m^{\prime}}\ .$ (5.9)
We observe that this algebra was defined in [24] in a slightly different way.
The KM and Virasoro algebra of the two-sphere admits a semi-direct structure $\widetilde{\mathfrak{g}}(\mathbb{S}^{2})\rtimes$Vir$(\mathbb{S}^{2})$ with action of $L_{\ell m}$ on $\mathcal{T}_{a\ell^{\prime}m^{\prime}}$:
$\big{[}L_{\ell m},\mathcal{T}_{a\ell^{\prime}m^{\prime}}\big{]}=-m^{\prime}c_{\ell,m,\ell^{\prime},m^{\prime}}^{\ell^{\prime\prime}}\mathcal{T}_{a\ell^{\prime\prime}m+m^{\prime}}\ ,$ (5.10)
and we have $J_{0}=-L_{00}$. The algebra $\widetilde{\mathfrak{g}}(\mathbb{S}^{2})\rtimes$Vir$(\mathbb{S}^{2})$ is thus defined by (5.5) and (5.9) and (5.10).
5.3 A KM algebra associated to $\mathbb{S}^{2}_{F}$
Consider now a ‘softened’ two-sphere $\mathbb{S}_{F}^{2}$ with a deformed metric $g_{MN}^{\mu}$ such that $\sqrt{\det{(g_{MN}^{\mu})}}=F(\theta,\psi)\sin\theta$. The scalar product on $L^{2}(\mathbb{S}_{F}^{2})$ reduces to
$(f,g)_{F}=\frac{1}{4\pi^{2}}\int\limits_{0}^{\pi}\text{d}\theta\int\limits_{0}^{2\pi}\text{d}\psi\sin\theta\;F(\theta,\psi)\;\overline{f(\theta,\psi)}g(\theta,\psi)\ .$
The results of Sections 2.2 and 2.3 enable us to define KM algebras associated to $\mathbb{S}_{F}^{2}$:
  1. 1.
    The Hilbert basis of $L^{2}(\mathbb{S}_{F}^{2})$ is
    $\mathcal{B}_{F}=\Big{\{}Y_{\ell m}^{F}(\theta,\psi)=\frac{1}{\sqrt{F\left(\theta,\psi\right)}}Y_{\ell m}\left(\theta,\psi\right)=P_{\ell m}^{~{}~{}\ell^{\prime}m^{\prime}}Y_{\ell^{\prime}m^{\prime}}(\theta,\psi)\ ,\ \ \ell\in\mathbb{N},-\ell\leq m\leq\ell\Big{\}},$
    where the coefficients $P_{\ell m}^{~{}~{}\ell^{\prime}m^{\prime}}$ are defined according to the general formula (2.26).
  2. 2.
    the Hermitean operator associated to $\mathbb{S}_{F}^{2}$ is $J_{0}$ itself.
  3. 3.
    the two-cocycle associated to $J_{0}$ is given by
    $\omega^{F}\left(T_{a_{1}\ell_{1}m_{1}}^{F},T_{a_{2}\ell_{2}m_{2}}^{F}\right)$ $=$ $P_{\ell_{1}m_{1}}^{~{}~{}~{}\ell_{1}^{\prime}m_{1}^{\prime}}P_{\ell_{2}m_{2}}^{~{}~{}~{}~{}~{}\ell_{2}^{\prime}m_{2}^{\prime}}\omega\left(T_{a_{1}\ell_{1}^{\prime}m_{1}^{\prime}},T_{a_{2}\ell_{2}^{\prime}m_{2}^{\prime}}\right)$
    $=$ $k\eta_{a_{1}a_{2}}\left(-1\right)^{m_{1}^{\prime}}m_{2}^{\prime}P_{\ell_{1}m_{1}}^{~{}~{}~{}\ell_{1}^{\prime}m_{1}^{\prime}}P_{\ell_{2}m_{2}}^{~{}~{}~{}~{}~{}\ell_{2}^{\prime}m_{2}^{\prime}}\delta_{\ell_{1}^{\prime}\ell_{2}^{\prime}}\delta_{m_{1}^{\prime},-m_{2}^{\prime}}.$
Thus from Sections 2.2 and 2.3, we define the analogue of the loop algebra $\mathfrak{g}(\mathbb{S}_{F}^{2})=\big{\{}T_{a\ell m}=T_{a}P_{\ell m}^{~{}~{}\ell^{\prime}m^{\prime}}Y_{\ell^{\prime}m^{\prime}}(\theta,\psi)\big{\}}$ and the corresponding centrally extended Lie algebra $\widetilde{\mathfrak{g}}(\mathbb{S}_{F}^{2})$,
$\widetilde{\mathfrak{g}}(\mathbb{S}_{F}^{2})=\Big{\{}\mathcal{T}_{a\ell m}^{F},J_{0},k\ ,\ \ a=1,\cdots,d,\ell\in\mathbb{N},-\ell\leq m\leq\ell\Big{\}}.$
Using the relation
$\frac{1}{\sqrt{F(\theta,\psi)}}=\sum_{\ell\in\mathbb{N}}\sum\limits_{m=-\ell}^{\ell}F^{\ell m}Y_{\ell m}(\theta,\psi)$
the Lie brackets reduce to
$\big{[}\mathcal{T}_{a_{1}\ell_{1}m_{1}}^{F},\mathcal{T}_{a_{2}\ell_{2}m_{2}}^{F}\big{]}$ $=$ $if_{a_{1}a_{2}}{}^{a_{3}}c_{\ell_{1},m_{1},\ell_{2},m_{2}}^{F\ell}\mathcal{T}_{a_{3},\ell_{3},m_{1}+m_{2}}^{F}+(-1)^{m_{1}^{\prime}}km_{2}^{\prime}\eta_{a_{1}a_{2}}P_{\ell_{1}m_{1}}^{~{}~{}\ell_{1}^{\prime}m_{1}^{\prime}}P_{\ell_{2}m_{2}}^{~{}~{}\ell_{2}^{\prime}m_{2}^{\prime}}\delta_{\ell_{1}^{\prime},\ell_{2}^{\prime}}\delta_{m_{1}^{\prime},-m_{2}^{\prime}},$
$\big{[}J_{0},\mathcal{T}_{a\ell m}^{F}\big{]}$ $=$ $m^{\prime}P_{\ell m}^{~{}~{}\ell^{\prime}m^{\prime}}\mathcal{T}_{a\ell^{\prime}m^{\prime}}^{F}.$ (5.11)
where the ‘Clebsch-Gordan coefficients’ of $\mathbb{S}_{F}^{2}$ $c_{\ell m\ell^{\prime}m^{\prime}}^{F\ell^{\prime\prime}}$ are defined by an expression analogous to (2.25).
If we now define an analogue of Virasoro algebra on the ‘softened’ two-sphere $\mathbb{S}_{F}^{2}$, by arguments similar than in Section 3, one can show that the Virasoro algebra of the ‘softened’ two-sphere is isomorphic to the Virasoro algebra on the two-sphere.
5.3.1 $\widetilde{\mathbb{S}}^{2}$ and its physical applications
The squashed two-sphere $\mathbb{S}_{b}^{2}\equiv\mathbb{\tilde{S}}^{2}$ is a ‘softened’ version of the usual two-sphere $\mathbb{S}^{2}$, where the geometry is compressed or stretched along one axis, breaking the spherical symmetry while maintaining some degree of residual symmetry, typically $U(1)$. The metric of the squashed two-sphere is often written as
$ds^{2}=d\theta^{2}+b^{2}\sin^{2}\theta\,d\phi^{2}$
where $b$ is the squashing parameter, $\theta\in[0,\pi]$ and $\phi\in[0,2\pi)$ are the usual spherical coordinates. For $b=1$, this reduces to the standard metric on $\mathbb{S}^{2}$, but when $b\neq 1$, the symmetry is reduced to $U(1)$, representing squashing along one of the axes. This squashed geometry has important applications in theoretical physics, especially in gauge theory [66], string theory [67], and supersymmetric field theory [68].
[…]
In string theory, $\mathbb{\tilde{S}}^{2}$ appears in the context of flux compactifications and as an internal space in string sigma models. For instance, in compactifications of string theory on non-trivial backgrounds, $\mathbb{\tilde{S}}^{2}$ provides a natural compactification space with reduced symmetry that still preserves some supersymmetry. This compactification can lead to interesting low-energy effective theories where the squashing parameter $b$ controls the amount of symmetry breaking. Additionally, in string sigma models $\mathbb{\tilde{S}}^{2}$ provides a target space for two-dimensional field theories that describe strings propagating on deformed geometries. The deformation of the target space modifies the spectrum of the theory, leading to shifts in the masses of excitations and affecting the dynamics of the system [67].
Last but not least, within the AdS/CFT correspondence, as mentioned above squashed spheres, thus including $\mathbb{\tilde{S}}^{2}$, arise in the context of holographic dualities. In particular, squashed spheres appear as internal compactification spaces in AdS spacetimes, where the dual field theory resides on the boundary of the AdS space. The squashing of the internal sphere leads to deformations of the dual field theory, affecting its operator spectrum and the correlation functions. For example, $\mathbb{\tilde{S}}^{2}$ can modify the dual conformal field theory by breaking certain symmetries while preserving others, thereby providing a useful tool to study symmetry-breaking phenomena in the holographic setting [68].
The invariant measure on the squashed sphere reads
$\int\limits_{{\mathbb{S}^{2}}}\text{d}\mu(\widetilde{\mathbb{S}^{2}})=\frac{b}{4\pi}\int\limits_{0}^{\pi}\sin\theta\text{d}\theta\int\limits_{0}^{2\pi}\text{d}\psi\ .$ (5.12)
This means that
$\mathcal{B}_{\widetilde{\mathbb{S}}^{2}}=\Big{\{}\widetilde{Y}_{\ell m}=\frac{1}{\sqrt{b}}Y_{\ell m}\ ,\ \ \ell\in\mathbb{N}\ ,\ \ -\ell\leq m\leq\ell\Big{\}}$
with $Y_{\ell m}$ corresponding to the matrix elements of $\mathbb{S}^{2}$ (see e.g. (5.2)) constitutes a Hilbert basis of the squashed sphere:
$(\widetilde{Y}_{\ell m},\widetilde{Y}_{\ell^{\prime}m^{\prime}})=\delta_{\ell\ell^{\prime}}\delta_{mm^{\prime}}\ .$
Note that, as mentioned in Section 4.2, $\widetilde{Y}_{\ell m}$ are in a representation of SO$(3)$, but the isometry of the squashed sphere reduces to U$(1)$. From Section 5.1, the description of the product $\widetilde{Y}_{\ell m}(\theta,\psi)\widetilde{Y}_{\ell^{\prime}m^{\prime}}(\theta,\psi)$ is straightforward. The KM algebra associated to the squashed sphere $\widetilde{\mathfrak{g}}(\widetilde{\mathbb{S}}^{2})$ follows directly, and it is isomorphic to the KM algebra associated to the two-sphere. Since the usual sphere is obtained taking the limit $b\to 1$, considering this limit we recover an isomorphic realization of the KM algebra of $\mathbb{S}^{2}$ from the KM algebra of the squashed sphere.
6 Conclusions
In this paper, we have proposed another generalization of infinite-dimensional (and infinite-rank) generalized KM algebras $\mathfrak{g}\left(\mathcal{M}\right)$ (and their central extensions $\mathfrak{\tilde{g}}\left(\mathcal{M}\right)$) introduce in [20, 21]. Namely, we have set $\mathcal{M}=G_{c}$, and considered the compact group manifold $G_{c}$ to be ‘deformed’ into a so-called ‘soft’ group manifold $G_{c}^{\mu}$, locally diffeomorphic to $G_{c}$ itself. This ‘softening’ procedure lies at the core of the group-geometric approach to (super)gravity and superstring theories, and it was extensively studied from the 1970s onwards by Tullio Regge and his research group, being further developed later by Riccardo D’Auria and Pietro Fré in Turin [32]. In this context, the ‘softening’ corresponds to deform the original rigid group manifold structure of $G_{c}$, where the left- and right- invariant vector fields and one-forms locally have a fixed coordinate dependence, and where the Riemannian geometry is (locally) fixed in terms of the structure tensor of (the Lie algebra $\mathfrak{g}_{c}$ of) $G_{c}$. In this context, it is worth mentioning that some work of Castellani [28, 69], in which he proposed some pioneering generalizations of KM algebras, shares some intriguing structural properties with the ansatz considered in this paper. The exact implications deriving from the comparison of both approaches are not yet fully explored, and their detailed analysis is left for a future investigation.
Thus, the algebraic generalization achieved in this paper should be regarded as a relevant structural step towards the application of the generalized KM algebras in the context of (super)gravity and (super)string theories. As explicit examples, we have considered the ‘softening’ of the one-sphere $\mathbb{S}^{1}$, of the two-sphere $\mathbb{S}^{2}$, and of the three-sphere $\mathbb{S}^{3}$ (which in particular include the squashed three-sphere $\widetilde{\mathbb{S}}^{3}$, also named Berger three-sphere). While the generalized KM algebra associated to the ‘softened’ circle $\mathbb{S}_{F}^{1}$ is trivially isomorphic to its undeformed analogue and thus deprived of interest, the ‘softening’ of $\mathbb{S}^{2}$ and $\mathbb{S}^{3}$ yields to non-trivial results, which would be very interesting to apply to the broad range of contexts (briefly reviewed within our treatment) in which their undeformed counterparts play an important role. In this framework, by developing the suggestion made in the conclusion of [28] (in which a generalization of general relativity for closed strings was formulated), the generalized KM algebras associated to $\mathbb{S}^{2}$ and $\mathbb{S}^{3}$ would be expected to yield an extension of general relativity for closed 2- and 3- branes, respectively; the physical meaning of the corresponding ‘softenings’ remains an intriguing venue for further future research.
Of course, along the lines of research pertaining to the present paper, further possible developments would consist into applying the ‘softening’ procedure to non-compact Lie group manifolds $G_{nc}$ (or cosets thereof) on which the generalized KM algebras introduced in [29, 30] are based; consequently, $G_{nc}$ would be ‘softened’ into the manifold $G_{nc}^{\mu}$. Indeed, it should be here recalled that the prototypical example of ‘softening’ is based on $G_{nc}/H=ISO(3,1)/SO(3,1)$, thus realizing the four-dimensional space-time manifold in which dynamical fields are defined through an ‘horizontalization’ procedure (cf. e.g. [32]). Therefore, the infinite-dimensional algebras resulting from this further step of algebraic generalization would be based on the non-compact ‘soft’ group manifold $G_{nc}^{\mu}$, and they would be potentially relevant for the formulation of gravitational theories, with or without underlying supersymmetry; in this respect, it would be interesting to investigate possible relations with the results of the recent paper [70]. We leave this intriguing venue of investigation for further future research.
Acknowledgments
The work of RCS has been supported by the Agencia Estatal de Investigación (Spain) under the grant PID2023-148373NB-I00 funded by MCIN/AEI/10.13039/501100011033/FEDER, UE. This article is based upon work from COST Action CaLISTA CA21109 supported by COST (European Cooperation in Science and Technology).
Appendix A Maurer-Cartan one-forms for $SU(2)$
For any Lie group (manifold) $G$, the $G$-left-invariance of the ‘left’ $\mathfrak{g}$-valued Maurer-Cartan one-form (where $\mathfrak{g}=Lie\left(G\right)$ is the Lie algebra of $G$) is immediate (see e.g. [71]). Indeed, the ‘left’ Maurer-Cartan one-form $\lambda$ is defined as
$\lambda:=g^{-1}dg,$
whereas the left multiplication by an element $h\in G$, denoted as $L_{h}$, acts on a group element $g\in G$ as $L_{h}\left(g\right):=hg$. Thus, for any $g\in G$, the differential of the map $L_{h}$ on the group manifold $G$ is
$d\left(L_{h}\left(g\right)\right)=d\left(hg\right)=hdg.$
The invariance of $L_{h}$-transformed one-forms $\lambda^{\prime}:=\left(L_{h}\left(g\right)\right)^{-1}d\left(L_{h}\left(g\right)\right)$ follows at once from the action, as
$\lambda^{\prime}:=\left(L_{h}\left(g\right)\right)^{-1}d\left(L_{h}\left(g\right)\right)=\left(hg\right)^{-1}hdg=g^{-1}h^{-1}hdg=g^{-1}dg=\lambda.~{}~{}\square$
Analogously, one can prove the $G$-right-invariance of the ‘right’ $\mathfrak{g}$-valued Maurer-Cartan one-form, defined as
$\omega:=\left(dg\right)g^{-1}.$
The right multiplication by an element $h\in G$, denoted as $R_{h}$, acts on a group element $g\in G$ as $R_{h}\left(g\right):=gh$. Thus, for any $g\in G$, the differential of the map $R_{h}$ on the group manifold $G$ is
$d\left(R_{h}\left(g\right)\right)=d\left(gh\right)=\left(dg\right)h.$
Again, the $R_{h}$-transformed Maurer-Cartan one-form $\omega^{\prime}:=\left(d\left(R_{h}\left(g\right)\right)\right)\left(R_{h}\left(g\right)\right)^{-1}$ coincides with itself :
$\omega^{\prime}:=\left(d\left(R_{h}\left(g\right)\right)\right)\left(R_{h}\left(g\right)\right)^{-1}=\left(dg\right)h\left(gh\right)^{-1}=\left(dg\right)hh^{-1}g^{-1}=\left(dg\right)g^{-1}=\omega.~{}~{}\square$
The $SU(2)$ group manifold can be parameterised using three Euler angles $\varphi$, $\theta$, and $\psi$, which correspond to a sequence of rotations about specific axes. As $SU(2)$ is the double cover of the rotation group $SO(3)$, each rotation in $SO(3)$ corresponds to two points in $SU(2)$, providing a complete representation of all possible orientations. The three Euler angles allow us to express any element $g\in SU(2)$ as a product of rotations as follows:
$g(\varphi,\theta,\psi)=e^{i\psi\sigma_{3}/2}e^{i\theta\sigma_{2}/2}e^{i\varphi\sigma_{3}/2}$
where $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{3}$ are the Pauli matrices:
$\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\sigma_{2}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\quad\sigma_{3}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
The range of the angles are as follows : $\psi\in[0,4\pi)$ (initial rotation about the $z$-axis); $\theta\in[0,\pi]$ (rotation about the $y$-axis); $\varphi\in[0,2\pi)$ (final rotation about the $z$-axis). With these parameters, the generic element $g(\psi,\theta,\varphi)\in SU(2)$ can be written explicitly as:
$g(\psi,\theta,\varphi)=\begin{pmatrix}e^{i(\psi+\varphi)/2}\cos(\theta/2)&-e^{i(\psi-\varphi)/2}\sin(\theta/2)\\ e^{-i(\psi-\varphi)/2}\sin(\theta/2)&e^{-i(\psi+\varphi)/2}\cos(\theta/2)\end{pmatrix}\in SU(2)\simeq\mathbb{S}^{3},$
(to be compared with (4.1)). This parameterisation provides a complete description of any element in $SU(2)$ through the three Euler angles $\varphi$, $\theta$, and $\psi$. The double covering property of $SU(2)$ over $SO(3)$ is reflected in the range of $\psi$ extending to $4\pi$, ensuring that each rotation in $SO(3)$ corresponds to two points in $SU(2)$.
Using this parameterisation, one can easily compute the right-invariant Maurer-Cartan one-forms $\omega$’s (4.1). These are obtained by taking the differential of the group element and transforming it by the inverse of the group element on the right. Specifically, if $g(\psi,\theta,\varphi)$ is a generic element of $SU(2)$, then the right-invariant Maurer-Cartan one-form $\omega$ is given by:
$\omega=(dg)g^{-1}$
where $\omega$ is an $\mathfrak{su}(2)$-valued one-form. In this case, we can expand $\omega$ as a linear combination of the Pauli matrices $\sigma_{i}$, which form a basis for $\mathfrak{su}(2)=Lie\left(SU(2)\right)$. Thus, one obtains
$\omega=\frac{i}{2}\left(\omega_{1}\sigma_{1}+\omega_{2}\sigma_{2}+\omega_{3}\sigma_{3}\right)$
where $\omega_{1}$, $\omega_{2}$, and $\omega_{3}$ are given by (4.1). Note that $\omega_{3}$ is also invariant under $L_{0}$.
Analogously, one can compute the left-invariant Maurer-Cartan one-forms $\lambda$’s (4.1), defined as
$\lambda=g^{-1}dg=\frac{i}{2}\left(\lambda_{1}\sigma_{1}+\lambda_{2}\sigma_{2}+\lambda_{3}\sigma_{3}\right).$
Note that $\lambda_{3}$ is also invariant under $R_{0}$. An equivalent way to obtain the one-forms $\lambda$’s is to observe that
$\left.R_{\pm}\right|_{\varphi\leftrightarrow\psi}=-L_{\pm},~{}\left.R_{0}\right|_{\varphi\leftrightarrow\psi}=L_{0},$
and thus ($\forall i=1,2,3$)
$\lambda_{i}=\left.\omega_{i}\right|_{\varphi\leftrightarrow\psi},$
as it follows immediately from (4.1) and (4.1).